### v1

# RePO: ReLU-based Preference Optimization

Abstract
Aligning large language models (LLMs) with human preferences is critical for real-world deployment, yet existing methods like RLHF face computational and stability challenges. While DPO establishes an offline paradigm with single hyperparameter $\beta$, subsequent methods like SimPO reintroduce complexity through dual parameters ($\beta$, $\gamma$). We propose ReLU-based Preference Optimization (RePO), a streamlined algorithm that eliminates $\beta$ via two advances: (1) retaining SimPO’s reference-free margins but removing $\beta$ through gradient analysis, and (2) adopting a ReLU-based max-margin loss that naturally filters trivial pairs. Theoretically, RePO is characterized as SimPO’s limiting case ($\beta\to\infty$), where the logistic weighting collapses to binary thresholding, forming a convex envelope of the 0-1 loss. Empirical results on AlpacaEval 2 and Arena-Hard show that RePO outperforms DPO and SimPO across multiple base models, requiring only one hyperparameter to tune.
LLM Alignment, RLHF, Data Selection
1 Introduction
Aligning large language models (LLMs) with human preferences is essential for their effective deployment in real-world applications (Anil et al., 2023; Touvron et al., 2023; OpenAI, 2023; Bubeck et al., 2023), ensuring that their outputs are beneficial, accurate, and adhere to human values (Ouyang et al., 2022) while minimizing potential safety risks. A primary approach for achieving this alignment is Reinforcement Learning from Human Feedback (RLHF). It typically entails a multi-stage process: initially, a reward model is trained using ranked preference data; subsequently, the LLM is optimized with the guidance of reward model, using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017). Despite its effectiveness, RLHF faces significant challenges, such as high computational costs and training instability (Rafailov et al., 2023; Zhao et al., 2023).
Recent research has been exploring simpler offline alternatives. Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one such approach. It reparameterizes the reward function in RLHF, bypassing the explicit learning of the reward model, and thus enabling direct training of LLMs with human preference data. Formally, the DPO objective is a logistic-log loss, which applies a log-sigmoid activation to the reward margin between the preferred and less-preferred responses to the same input. Here, the reward is derived from comparing the policy model of the LLM with a reference model. Building on this, follow-up studies, especially SLiC-HF (Zhao et al., 2023) and SimPO (Meng et al., 2024), focus on further simplifications on the loss function and reward margins separately. Specifically, SLiC-HF replaces the logistic-log loss in DPO with a hinge loss, achieving a faster-decaying tail. SimPO, on the other hand, introduces reference-free reward margins with logistic-log loss, offering a more streamlined and efficient approach. While advancing the field and achieving leading performance, they introduce additional complexity on hyperparameter tuning. SimPO inherits DPO’s reward margin strength controller $\beta$ and adds a target reward margin $\gamma$. Similarly, SLiC-HF utilizes $\gamma$, but also introduces a regularization weight $\lambda$. The necessity of dual-tuning these hyperparameters poses significant challenges in model training and deployment. This leads to a natural question:
Can we develop a simpler offline preference optimization algorithm?
Figure 1: RePO, SimPO, and DPO primarily differ in their loss functions, as highlighted in the shaded box. RePO either outperforms or achieves comparable performance to DPO and SimPO across various settings on AlpacaEval 2, with only a single hyperparameter $\gamma$. “IT” denotes pre-trained instruction-tuned models. e.g., “Gemma2-IT 9B” refers to Gemma2-9B (instruct setup).
In this paper, we propose a simple idea: simultaneously simplifying the log-sigmoid activation and reward margins, to achieve comparable or even better performance. For reward margins, we adopt the reference-free approach of SimPO but eliminate the hyperparameter $\beta$. For the log-sigmoid activation, we employ ReLU to the reference-free reward margins, following a max-margin (support vector machine) paradigm (Boser et al., 1992; Cortes & Vapnik, 1995), similar to SLiC-HF. Consequently, our method can be interpreted as “SimPO without $\beta$” and “SLiC-HF without $\lambda$”. Interestingly, it bridges these methods by omitting a core component from each while maintaining competitive performance, demonstrating the effectiveness of this streamlined approach. We term this ReLU-based Preference Optimization RePO.
Theoretically, we demonstrate that RePO emerges as the natural endpoint of SimPO’s architectural evolution: when $\beta\to\infty$, SimPO’s logistic weighting collapses to RePO’s binary thresholding (Lemma 3.1). This asymptotic behavior motivates our key theoretical insight — the ReLU loss in RePO is precisely the convex envelope of the 0-1 loss (Theorem 4.2), inheriting its global optimality while enabling tractable gradient-based optimization. Practically, this formulation provides the following advantages:
  • Simplified Hyperparameter Tuning: RePO removes the hyperparameter $\beta$, requiring only the tuning of $\gamma$ within the range $[0,1]$. Empirical results show that RePO can achieve performance comparable to SimPO on three models (Llama, Mistral, and Gemma) even with a fixed $\gamma=0.5$, significantly reducing the complexity of hyperparameter tuning.
  • Effective Data Filtering: By using ReLU instead of a continuous sigmoid function, RePO interprets the target reward margin $\gamma$ as a strict “cutoff point”. This mechanism zeroes out gradients for preference pairs whose reward margin exceeds $\gamma$ (well-separated pairs), while assigning unit weight to less-separated pairs. This selective approach minimizes overfitting by focusing optimization on challenging data points.
  • Controllable Over-Optimization: The “cutoff point” $\gamma$ controls the mean of reward margin within a batch, representing a new metric for evaluating over-optimization. Experimental results suggest that this metric correlates with model behavior and can replace KL divergence as a simpler alternative, potentially controlling optimization without requiring a reference model.
While RePO effectively filters out well-separated pairs by zeroing their gradients, it assigns a unit gradient weight to all less-separated pairs. This implicit uniform weighting overlooks the varying importance of these pairs, potentially limiting its ability to prioritize more informative examples. To address this limitation, we extend the concept of RePO within SimPO and propose a new method, RePO++, by leveraging SimPO’s logistic-log loss to differentiate the importance of less-separated pairs. This approach assigns greater weights to less-separated pairs, while continuing to filter out well-separated pairs. Notably, only a single line of code is required to seamlessly integrate RePO into SimPO, resulting in RePO++. In a nutshell, our main contribution is ReLU-based Preference Optimization (RePO), a simple but effective reference-free algorithm for training language models from preferences. With its straightforward motivation, RePO is grounded in a principled theoretical framework and consistently outperforms DPO and SimPO across three base model settings (Mistral2-7B, Llama3-8B, and Gemma2-9B) on AlpacaEval 2 (cf. Figure 1). We hope our exploration will motivate people to rethink the fundamental roles of loss function design for preference optimization.
2 Preliminaries
Offline Alignment. In the offline alignment setting, we utilize a dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$, where each entry consists of a prompt $x$ accompanied by two responses: $y_{w}$ is the preferred (winning) response, and $y_{l}$ is the less preferred (losing) response, both generated by a reference policy $\pi_{\text{ref}}$. The primary challenge is to find a latent reward function $r^{*}(x,y)$ that governs these preference relationships. Although $r^{*}(x,y)$ is not directly observable, the Bradley-Terry (BT) model (Bradley & Terry, 1952) provides a probabilistic framework for modeling the probability of $y_{w}$ over $y_{l}$ given $x$:
$\mathbb{P}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r^{*}(x,y_{w}))}{\exp(r^{*}(x,y_{w}))+\exp(r^{*}(x,y_{l}))}.$ (1)
Here, $r^{*}(x,y)$ represents the latent reward for a response $y$ in the context of prompt $x$. The aim of offline alignment is to develop a policy $\pi_{\theta}$ capable of accurately estimating this latent reward function using the dataset $\mathcal{D}$.
Reinforcement Learning from Human Feedback (RLHF). Traditional offline alignment methods often leverage reinforcement learning augmented with a KL-divergence regularization term in the reward objective (Bai et al., 2022; Ziegler et al., 2019; Ouyang et al., 2022). The optimization objective, parameterized by a regularization parameter $\beta>0$, is expressed as:
$\mathop{\max}_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y\mid x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)),$ (2)
where $r_{\phi}(x,y)$ represents the reward function trained using the BT model on the preference data, $\pi_{\theta}$ is the policy being optimized, and $\pi_{\text{ref}}$ is the fixed reference policy, typically derived from supervised fine-tuning. The KL-divergence term ensures that $\pi_{\theta}$ remains close to $\pi_{\text{ref}}$, mitigating risks of excessive divergence during training.
Directed Preference Optimization (DPO). DPO (Rafailov et al., 2023) stands out as a leading method for offline preference optimization by eliminating the need for an explicit reward model. Instead, it reformulates the reward function $r(x,y)$ as a closed-form expression based on policy ratios:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x),$ (3)
where $Z(x)$ is a partition function that does not depend on $y$. This leads to the DPO loss for a given triplet $(x,y_{w},y_{l})$ as:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}$ (4)
$\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right)\right],$
where $\sigma(\cdot)$ denotes the sigmoid function. This loss encourages the policy $\pi_{\theta}$ to prefer $y_{w}$ over $y_{l}$ in alignment with the reference policy.
Sequence Likelihood Calibration (SLiC-HF). SLiC-HF (Zhao et al., 2023) advances preference optimization with two key innovations: (1) it employs a sequence-level calibration loss that contrasts the log-probability difference between preferred and dispreferred responses using a margin $\gamma$, and (2) it integrates a regularization term to prevent divergence from the SFT policy, avoiding the need for an explicit KL penalty. The SLiC-HF loss function is defined as:
$\mathcal{L}_{\text{SLiC-HF}}(\pi_{\theta})=$ $\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\Big{[}\text{ReLU}\Big{(}-\Big{(}\log{\pi_{\theta}(y_{w}\mid x)}$ (5)
$-\log\pi_{\theta}(y_{l}\mid x)-\gamma\Big{)}\Big{)}-\lambda\log{\pi_{\theta}(y_{w}\mid x)}\Big{]}.$
Simple Preference Optimization (SimPO). SimPO (Meng et al., 2024) advances preference optimization with two key innovations: (1) it normalizes the reward by the length of the response, calculating the average log-probability per token for a response under the policy $\pi_{\theta}$, and (2) it incorporates a target reward margin $\gamma$ to ensure that the reward difference between the preferred and less preferred responses exceeds this margin. The SimPO loss function is defined as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}$ (6)
$\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|}-\gamma\right)\right)\right],$
where $|y|$ denotes the number of tokens in response $y$, ensuring length-aware scaling of rewards, and $\gamma$ is the predefined margin that enforces a minimum difference in rewards between $y_{w}$ and $y_{l}$. To align with subsequent discussions, we modify the original SimPO formulation by setting $\gamma$ to $\gamma/\beta$.
3 Analysis on RePO
In this section, we first present the expression for RePO and compare its performance with that of SimPO through experiments. We then analyze the advantages and key factors behind the success of RePO from the perspectives of gradient behavior, data filtering, and over-optimization.
3.1 RePO
Method Formulation. We propose RePO by simplifying the SimPO loss function through two key modifications: (i) removing the hyperparameter $\beta$, and (ii) replacing the log-sigmoid function with a ReLU activation. First, we define the implicit reward margin $M_{\theta}$ as:
$M_{\theta}=\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|},$ (7)
which measures the policy’s relative preference between responses. Using $M_{\theta}$, the proposed loss function is:
$\mathcal{L}_{\text{RePO}}(\pi_{\theta})=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\mathrm{ReLU}\left(-(M_{\theta}-\gamma)\right)\right],$ (8)
where $\gamma\in[0,1]$ is the sole hyperparameter representing the target reward margin.
Comparison between RePO and SimPO / SLiC-HF. As formalized in Equation 8, RePO simplifies the SimPO framework by replacing the log-sigmoid activation with a ReLU operator and removing the need to tune the hyperparameter $\beta$. Simultaneously, it eliminates the explicit regularization term in SLiC-HF, thereby removing the associated hyperparameter $\lambda$. This dual simplification makes RePO a unified specialization of both SimPO and SLiC-HF, reducing hyperparameter tuning complexity while maintaining competitive performance, as demonstrated in Section 3.2.
Training Dynamics Monitoring. We track optimization progress with two metrics:
  • Batch-level reward margin:
    $m_{\text{batch}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{B}}[M_{\theta}],$ (9)
    which reflects the average reward margin within each training batch $\mathcal{B}$.
  • Dataset-level reward margin:
    $m_{\mathcal{D}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}[M_{\theta}],$ (10)
    which measures the overall reward margin on the entire training set $\mathcal{D}$.
Here, $m_{\text{batch}}$ provides immediate feedback during each update, while $m_{\mathcal{D}}$ captures the global preference strength of the policy over the complete dataset.
Gradient Analysis. We analyze the gradient dynamics of RePO and SimPO to highlight their fundamental differences, the gradients are as follows:
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\beta\mathbb{E}_{\mathcal{D}}\left[s_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right],$ (11)
$\nabla_{\theta}\mathcal{L}_{\text{RePO}}(\pi_{\theta})=-\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}{(M_{\theta}<\gamma)}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right],$ (12)
where $s_{\theta}=\sigma(\beta(-M_{\theta}+\gamma))$ is SimPO’s sigmoid weighting function. The terms $\nabla_{\theta,y_{w}}=\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}\mid x)$ and $\nabla_{\theta,y_{l}}=\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}\mid x)$ correspond to the gradients that increase the probability of the “winning” response $y_{w}$ and decrease the probability of the “losing” response $y_{l}$, respectively. The scaling factor $\beta$ in Equation 11 linearly amplifies gradient magnitudes but does not alter the relative update directions in adaptive optimizers like Adam (Kingma & Ba, 2014), as the momentum terms automatically normalize scale variations. We therefore omit $\beta$ in Figure 2 for clearer visualization of the weighting function shapes.
RePO’s gradient (Equation 12) applies uniform updates to pairs with $M_{\theta}<\gamma$, while SimPO (Equation 11) employs $\beta$-scaled continuous weighting through $s_{\theta}$. Figure 2 visualizes this contrast, showing how RePO’s binary thresholding emerges as a limiting case of SimPO when $\beta\to\infty$.
Figure 2: Gradient weighting functions of SimPO ($s_{\theta}$) and RePO ($\mathbb{I}(M_{\theta}<\gamma)$). As $\beta\to\infty$, $s_{\theta}$ converges to the binary indicator (dashed line), establishing RePO as the limit case of SimPO.
Lemma 3.1 (Gradient Equivalence in the SimPO-to-RePO Limit).
Under the same $M_{\theta}$ and $\gamma$ definitions, the SimPO gradient converges pointwise to the RePO gradient as $\beta\to\infty$:
$\lim_{\beta\to\infty}\nabla_{\theta}\mathcal{L}_{\text{SimPO}}=\nabla_{\theta}\mathcal{L}_{\text{RePO}}.$ (13)
Sketch.
The convergence follows from the pointwise limit of the sigmoid weighting:
$\lim_{\beta\to\infty}s_{\theta}=\lim_{\beta\to\infty}\sigma(\beta(-M_{\theta}+\gamma))=\mathbb{I}(M_{\theta}<\gamma).$
Substituting this into Equation 11 yields Equation 12. ∎
Remark 3.2.
Please check Appendix for all proofs. Lemma 3.1 establishes RePO as the asymptotic limit of SimPO with large $\beta$. This explains why RePO achieves comparable performance to SimPO (Section 3.2) while eliminating $\beta$ tuning. The binary thresholding in RePO also induces implicit data filtering, which we analyze in Section 3.3.
3.2 Relationship between SimPO and RePO
The previous section analyzes the relationship between SimPO and RePO from the perspective of gradient behavior. In this section, we compare their performance from an empirical standpoint.
Experimental setup. We adopt the experimental setup used in SimPO (Meng et al., 2024), with Llama3-8B and Gemma2-9B as our base models. For comparison, we use pre-trained instruction-tuned models as reference policies. To ensure consistency, we employ the same training datasets as in SimPO: princeton-nlp/llama3-ultrafeedback-armorm [^1] https://huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm for Llama3-8B and princeton-nlp/gemma2-ultrafeedback-armorm [^2] https://huggingface.co/datasets/princeton-nlp/gemma2-ultrafeedback-armorm for Gemma2-9B. In all experiments involving SimPO, we set $\beta=10.0$ and $\gamma=0.4$ for Gemma2-9B and $\beta=10.0$ and $\gamma=0.3$ for Llama3-8B, unless stated otherwise.
Figure 3: Performance of SimPO with varying $\beta$ and RePO on AlpacaEval2 benchmark.
Evaluation benchmarks. We evaluate our models on two widely used open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024). These benchmarks assess the models’ conversational abilities across various queries. For AlpacaEval 2, we report both the length-controlled win rate (LC) and raw win rate (WR). For Arena-Hard, we report the win rate (WR).
Observation 1: Large $\beta$ enhances SimPO’s performance. We first experiment with SimPO by varying the parameter $\beta$ over the set {1.0, 3.0, 5.0, 10.0}, while keeping $\gamma$ fixed. As shown in Figure 3, increasing $\beta$ consistently improves SimPO’s performance,with convergence observed as $\beta$ grows larger
Observation 2: RePO achieves competitive performance. RePO achieves performance comparable to SimPO with a large $\beta$. As shown in Figure 3, RePO achieves win rates of 51.1% on Llama3-8B and 66.6% on Gemma2-9B, comparable to SimPO’s performance. This aligns with Lemma 3.1, which establishes that RePO can be interpreted as a limiting case of SimPO as $\beta\to\infty$.
Figure 4: Line plot illustrating the variation in RePO’s performance with different values of $\gamma$, measured by the AlpacaEval 2 LC Win Rate. The accompanying bar chart shows the mean implicit reward margin, $m_{\mathcal{D}}$, computed across all training pairs.
Observation 3: $\gamma$ is a crucial hyperparameter. In RePO, the gradients become zero when the implicit reward margin exceeds $\gamma$, indicating that the model $\pi_{\theta}$ is confident in its predictions for these data pairs. These pairs, considered well-separated, are excluded from updates, mitigating overfitting. Thus, $\gamma$ serves as a critical “cutoff point”. As shown in Figure 4, the performance of RePO is highly sensitive to $\gamma$. Increasing $\gamma$ initially improves performance, but beyond a certain threshold, performance degrades. Intuitively, A small $\gamma$ retains many zero-gradient data points, leading to under-filtering. Conversely, a large $\gamma$ causes many data points to have unit gradient weight, potentially leading to overfitting. Further discussion on $\gamma$’s role in data filtering is provided in subsequent sections.
Figure 5: Implicit reward margin $M_{\theta}$ distribution across training steps (total: 467) for RePO at $\gamma=0.4$. Dashed line: $\gamma=0.4$. Green: samples below $\gamma$ (gradient descent); gray: samples above $\gamma$ (zero gradient). Numbers: fraction of samples above $\gamma$.
3.3 Data Filtering by RePO
Building upon the preceding discussion, we observe that adjusting the parameter $\gamma$ defines a “cutoff point” that determines which sample gradients are set to zero, effectively implementing data filtering. In this section, we analyze the training process of RePO through the lens of data filtering.
Observation 4: $\gamma$ controls the distribution of implicit reward margins. We also report $m_{\mathcal{D}}$ (cf. Equation 10), the mean of implicit reward margins across all training pairs. As shown in the bar chart of Figure 4, as $\gamma$ increases, $m_{\mathcal{D}}$ rises. However, alignment performance w.r.t. LC improves initially but declines once $\gamma$ exceeds a certain threshold. This suggests that while a larger $\gamma$ increases the separation between training pairs, excessive separation can introduce the risk of reward hacking (Amodei et al., 2016), leading to a performance drop.
Observation 5: Less data, better alignment. Figure 5 illustrates the distribution of the implicit reward margin $M_{\theta}$ across different training stages, alongside the ratio of zero gradients (i.e., when $M_{\theta}$ exceeds $\gamma$). As training progresses, the model’s ability to discriminate between winning and losing samples improves, resulting in a steady increase in both the implicit reward margin and the ratio of filtered data. Notably, the filtered data ratio rises from 48% to 58% between steps 200 and 400. Despite using only half of the samples for gradient updates, the model achieves optimal performance, indicating that RePO effectively utilizes a smaller, more informative subset of training data.
3.4 Over-Optimization Analysis
When a larger $\gamma$ is selected, RePO permits most data to bypass the filter, resulting in a continuous increase in the reward margin between winning and losing samples. This phenomenon, a manifestation of reward exploitation or hacking (Amodei et al., 2016), is often referred to as “over-optimization” in the RLHF literature, and has been empirically investigated in both controlled experiments (Gao et al., 2023) and user studies (Dubois et al., 2023).
Model Over-Optimization: Building on Rafailov et al. (2024), we investigate over-optimization in RePO, by evaluating six different values of $\gamma$ ($0.0,0.2,0.4,0.6,0.8,1.0$), each corresponding to varying levels of data filtering. Across all cases, we observe a distinct hump-shaped performance pattern: while moderate filtering improves alignment, excessive filtering causes performance to degrade, highlighting the over-optimization effect.
Figure 6: Over-optimization results for RePO on Llama3-8B-IT and Gemma2-9B-IT. Metrics include model LC win rate on AlpacaEval2 and raw win rate on Arena-Hard. Fitted dotted curves apply scaling laws from Gao et al. (2023) to direct alignment, with GPT-4 win rates substituting gold reward model scores.
Scaling Law Fits. Previous work (Gao et al., 2023; Rafailov et al., 2024) has established scaling laws for reward model scores as a function of the KL divergence between the initial and optimized policies. In contrast, we eliminate the reference model and the associated computational cost of calculating KL divergence. Instead, we use the mean implicit reward margin during training as a proxy metric. The reward function $R(d)$ is given by:
$R(d)=d(\alpha-\beta\log d),$ (14)
where $\alpha$ and $\beta$ are constants dependent on the reward model’s dataset size and parameter count, and $d=m_{\text{batch}}$. Without training a proxy reward model, we substitute GPT-4 win rates over dataset completions for the gold reward. Interestingly, we find that this scaling law provides an accurate relationship between $d$ and win rates for RePO.
3.5 RePO++
Building on the previous discussion, we identify the advantages of RePO in finding a “cutoff point” to control over-optimization. However, a limitation persists: for cases where the implicit reward margin is smaller than $\gamma$, their gradient weights become unit, failing to differentiate between samples.
One potential solution is to apply a non-linear activation function to weight the gradient of each pair. Specifically, we extend RePO to DPO and SimPO, introducing a new method, RePO++.
$\mathcal{L}_{\text{RePO++}}(\pi_{\theta})=-\mathbb{E}_{\mathcal{D}}\left[\log\sigma\left(-\text{ReLU}\left(-\beta\left(M_{\theta}-\gamma\right)\right)\right)\right],$ (15)
What does the RePO++ update do? We again consider analyzing the gradient of the loss function RePO++ for an intuitive understanding. The gradient w.r.t. the parameters $\theta$ is given by:
$\nabla_{\theta}\mathcal{L}_{\mathrm{RePO++}}(\pi_{\theta})=-\beta\mathbb{E}_{\mathcal{D}}\left[s_{\theta}\cdot\mathbb{I}_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right],$ (16)
where $s_{\theta}=\sigma\left(\beta\left(-M_{\theta}+\gamma\right)\right)$ and $\mathbb{I}_{\theta}$ is an indicator function that is 1 if $M_{\theta}<\gamma$ and 0 otherwise. Intuitively, This gradient combines the strengths of SimPO and RePO: it scales updates by $s_{\theta}$ and filters them using $\mathbb{I}_{\theta}$, ensuring that the model focuses on less-separated pairs, with smaller separations receiving higher weights.
Adaptation with RePO++. The core contribution of RePO++ lies in leveraging ReLU to mitigate over-optimization while preserving the standard workflow of preference optimization. This makes RePO++ easily adaptable to existing DPO-like methods. For instance, as shown in Equation 15, replacing $M_{\theta}$ with $\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}$ seamlessly integrates RePO++ into DPO, forming a ReLU-enhanced version of DPO.
4 Theoretical Analysis of RePO
Next, we relate preference optimization to a supervised binary classification task and show how RePO leverages the convex envelope of the 0-1 loss to achieve both theoretical optimality and efficient optimization.
Following Tang et al. (2024), we reformulate reward modeling as a supervised binary classification task. Given data pairs $(z,l)$ with $z\in\mathbb{R}^{k}$ and labels $l\in\{-1,1\}$, the goal is to learn a predictor ^$\hat{\ell}(z)\in\mathbb{R}$ such that $\mathrm{sign}(\hat{\ell}(z))$ matches $l$. The classification accuracy is given by: $\frac{1}{2}\mathbb{E}[\mathrm{sign}(\hat{\ell}(z)\cdot l)]+\frac{1}{2},$ and the corresponding loss function is:
$\mathcal{L}_{0\text{-}1}(\hat{\ell}):=\mathbb{E}\left[1-\mathrm{sign}(\hat{\ell}(z)\cdot l)\right].$
The above loss, known as the 0-1 loss is non-convex. We therefore approximate it using convex functions $f:\mathbb{R}\to\mathbb{R}$:
$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(\hat{\ell}(z)\cdot l)\right].$
In the context of reward modeling, given response pairs $(y_{w},y_{l})$ with preference $y_{w}\succ y_{l}$, we construct binary classification samples with $\ell=1$. For a pointwise reward model $r_{\phi}:\mathcal{Y}\to\mathbb{R}$, the natural parameterization is: ^$\hat{\ell}(y_{w},y_{l})=r_{\phi}(y_{w})-r_{\phi}(y_{l}),$ leading to the loss:
$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(r_{\phi}(y_{w})-r_{\phi}(y_{l}))\right].$
Different choices of $f(x)$ correspond to distinct loss functions. Specifically, when $f(x)=\mathbb{I}(x<0)$, the 0-1 loss is obtained; when $f(x)=-\log\sigma(x)$, we obtain the logistic loss used in SimPO; and when $f(x)=\text{ReLU}(-x)$, it corresponds to the loss function used in RePO.
Definition 4.1.
Let $D\subseteq\mathbb{R}$ be a closed convex set. The convex envelope of $\mathcal{L}_{0\text{-}1}$ over $D$ is:
$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)$
::=:= $\sup\left\{h(x)\mid h\text{ is convex},\,h\leq\mathcal{L}_{0\text{-}1}\,\forall x\in D\right\}.$ (17)
Theorem 4.2 (Convex Envelope of 0-1 Loss).
For $D=[-a,b]$ with $a,b>0$, the convex envelope of $\mathcal{L}_{0\text{-}1}(x)=\mathbb{I}(x<0)$ is:
$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)=\frac{1}{a}\mathrm{ReLU}(-x).$
Corollary 4.3 (Optimality Preservation).
Let $D\subseteq\mathbb{R}$ be convex. Then:
  1. 1.
    For binary classification with loss $\mathcal{L}_{0\text{-}1}$:
    $\arg\min_{\hat{\ell}}\mathcal{L}_{0\text{-}1}(\hat{\ell})=\arg\min_{\hat{\ell}}\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(\hat{\ell}).$
  2. 2.
    When $D=[-a,b]$:
    $\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x).$
Remark 4.4.
Corollary 4.3 establishes: (i) equivalence between minimizing the 0-1 loss and its convex envelope, (ii) exact optimality preservation for the ReLU surrogate. This guarantees gradient-based optimization converges to solutions matching the theoretical optimum of the intractable 0-1 loss.
Corollary 4.5 (Logistic Loss is Not the Convex Envelope of 0-1 Loss).
The logistic loss $f_{\mathrm{log-sigmoid}}(x)=-\log\sigma(x)$ is not the convex envelope of $\mathcal{L}_{0\text{-}1}$.
Remark 4.6.
Unlike SimPO’s logistic loss (Corollary 4.5), the ReLU-based surrogate inherits the 0-1 loss’s global optimality. This explains RePO’s robustness to suboptimal local minima during gradient-based training.
5 Experiments
In this section, we present our main experimental findings, demonstrating that RePO consistently achieves better results than existing preference optimization methods across multiple benchmarks.
5.1 Experimental Setup
The main experimental setup has been described in Section 3.2. We increase a model: Mistral2-7B (Jiang et al., 2023). Additionally, the v0.2 Llama3-Instruct setup uses RLHFlow/ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024b) as the reward model for ranking generated data, significantly enhancing performance. As for baselines, we compare RePO with several state-of-the-art preference optimization methods: DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), IPO (Azar et al., 2023), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and R-DPO (Park et al., 2024). We also include the SFT model as a baseline. We thoroughly tune the hyperparameters for each baseline and report the best performance. Further details can be found in Appendix C.1. We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks (Beeching et al., 2023), with additional details in in Appendix C.2. The code is available at https://github.com/junkangwu/RePO.
Table 1: AlpacaEval 2 (AE2), Arena-Hard (AH) results across four settings. “WR” denotes the raw win rate,“LC” the length-controlled win rate. The best results are highlighted in bold, while the second-best are underlined.
Method Llama3-Instruct (8B) Mistral-Instruct (7B) Llama3-Instruct v0.2 (8B) Gemma2-Instruct (9B)
AE 2 AH AE 2 AH AE 2 AH AE 2 AH
LC (%) WR (%) WR (%) LC (%) WR (%) WR (%) LC (%) WR (%) WR (%) LC (%) WR (%) WR (%)
SFT 24.0 23.6 22.4 19.0 15.4 12.9 24.0 23.6 22.4 48.7 36.5 42.1
SLiC-HF 26.9 27.5 26.2 24.1 24.6 18.9 33.9 32.5 29.3 65.1 60.5 53.7
DPO 40.2 38.1 31.2 20.3 17.9 13.4 48.2 47.5 35.2 70.4 66.9 58.8
IPO 35.9 34.4 30.2 22.3 18.6 16.2 40.6 39.6 34.9 62.6 58.4 53.5
CPO 29.6 34.4 29.4 26.2 31.7 23.8 36.5 40.8 34.2 56.4 53.4 55.2
KTO 38.3 34.1 30.3 19.4 20.3 16.8 41.4 36.4 28.9 61.7 55.5 53.8
ORPO 31.6 29.8 26.3 24.0 23.0 18.6 36.5 33.1 30.4 56.2 46.7 46.2
R-DPO 40.3 37.3 32.9 21.4 22.2 13.8 51.6 50.7 35.0 68.3 66.9 57.9
SimPO 43.8 38.0 32.6 30.2 32.1 20.1 55.6 49.6 33.6 72.4 65.0 57.8
RePO 46.7 41.1 33.3 30.4 33.6 20.3 57.7 51.1 35.2 73.6 66.6 59.1
Table 2: Performance improvements of RePO and RePO++ over DPO and SimPO. Results are present on AlpacaEval 2 (AE 2) and Arena-Hard (AH) with LC (%) and WR (%). Red numbers indicate relative improvements.
Method Llama3-Instruct v0.2 (8B) Gemma2-Instruct (9B)
AE 2 AH AE 2 AH
LC (%) WR (%) WR (%) LC (%) WR (%) WR (%)
DPO 48.2 47.5 35.2 70.4 66.9 58.8
w. RePO $50.3^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.4\%}$ $51.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.1\%}$ $38.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+8.5\%}$ $73.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.8\%}$ $71.0^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.1\%}$ $64.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.2\%}$
w. RePO++ $50.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+5.4\%}$ $52.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.9\%}$ $37.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+5.7\%}$ $71.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.0\%}$ $69.5^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.9\%}$ $65.7^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+11.7\%}$
SimPO $55.6$ $49.6$ $33.6$ $72.4$ $65.0$ $57.8$
w. RePO $57.7^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.8\%}$ $51.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.0\%}$ $35.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.8\%}$ $73.6^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+1.7\%}$ $66.6^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.5\%}$ $59.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.2\%}$
w. RePO++ $56.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+0.9\%}$ $50.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+1.0\%}$ $35.9^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.8\%}$ $74.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.3\%}$ $66.5^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.3\%}$ $59.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.5\%}$
5.2 Main Results
RePO consistently and significantly outperforms existing preference optimization methods. As shown in Table 1, all preference optimization methods outperform the SFT model. Nevertheless, RePO — despite its simplicity — achieves the strongest overall performance across all benchmarks and settings. The consistent gains highlight its robustness and efficiency. Specifically, RePO surpasses the best-performing baseline by 0.2 to 2.8 points on the AlpacaEval 2 LC win rate under various configurations. On Arena-Hard, RePO generally performs better than competing methods, although CPO achieves slightly higher scores at times. We observe that CPO tends to produce responses that are approximately 50% longer. Because Arena-Hard does not penalize response length, longer outputs may occasionally yield higher evaluations.
5.3 Adaptations of RePO
RePO is a versatile approach that can be adapted to different optimization settings. In this work, we explore two adaptations: RePO and RePO++, applied to both DPO and SimPO. The detailed formulations are provided in Section 3.5.
As shown in Table 2, integrating RePO with DPO and SimPO consistently enhances performance. Notably, RePO++ achieves superior results in most scenarios, which we attribute to its combination of ReLU activation and the original weighting function $s_{\theta}$. This design effectively mitigates overoptimization while preserving the benefits of the original scheme. Interestingly, applying RePO and RePO++ to DPO yields particularly significant improvements (5%–12%), achieving a high score of 65.7 on the Arena-Hard benchmark.
5.4 Dynamic Margin Scheduling
To further investigate the impact of the target reward margin $\gamma$ , we conduct an ablation study using dynamic values for $\gamma$ Specifically, we implement two strategies: (i) $\gamma$ increases from small to large, and (ii) $\gamma$ decreases from large to small. Using RePO, we control the initial and final values of $\gamma$ , adjusting it uniformly over the course of one training epoch. As shown in Figure 7, both excessively small and large values of $\gamma$ lead to suboptimal performance (e.g., , $1.0\rightarrow 0.8$ , $0.0\rightarrow 0.2$ ). In contrast, starting with a moderately large value of $\gamma$ and gradually decreasing it during training (e.g., , $1.0\rightarrow 0.2$ ) results in improved model performance. We interpret this behavior as follows: in the early stages of training, the model is underfitting and thus benefits from stronger updates. In later stages, reducing $\gamma$ helps mitigate overfitting, effectively addressing this issue.
Figure 7: Performance of RePO with dynamic $\gamma$. The dashed line represents RePO with a fixed $\gamma$.
6 Discussion
Conclusion. In this work, we propose RePO, a simple preference optimization by replacing logistic weighting with a ReLU-based max-margin loss, eliminating the need for hyperparameter $\beta$. Theoretically, RePO is characterized as SimPO’s $\beta\rightarrow\infty$ limit, collapsing to binary thresholding for data filtering. Empirically, it matches or outperforms DPO and SimPO on AlpacaEval 2 and Arena-Hard with only a single tunable parameter ($\gamma$). Its gradient dynamics prioritize challenging samples, mitigating over-optimization while maintaining training stability. This work establishes RePO as a robust, hyperparameter-efficient alternative for offline alignment.
Limitations and Future Work. One limitation of our approach is that it currently operates in an offline setting, and future research will focus on extending it to online reinforcement learning frameworks. Integrating repo with online RL techniques could enhance real-time adaptability and scalability. Another direction for future work is to explore how the cutoff point strategy can be further refined to sustain performance improvements in self-play scenarios (Chen et al., 2024b), particularly in highly dynamic environments.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Related Works
Reinforcement learning from human feedback. RLHF is a technique designed to align large language models with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Azar et al., 2023). Traditional RLHF is typically structured in three stages: supervised fine-tuning (Zhou et al., 2023; Taori et al., 2023; Geng et al., 2023; Conover et al., 2023; Köpf et al., 2023; Ding et al., 2023), reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024a; Lightman et al., 2023; Havrilla et al., 2024; Lambert et al., 2024), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). In the third stage, Proximal Policy Optimization (PPO) is widely adopted. In contrast, RLOO (Ahmadian et al., 2024) reduces the GPU memory footprint of RLHF by eliminating the Critic model and leverages a Leave-One-Out strategy to achieve superior performance. GRPO (Shao et al., 2024), another variant of PPO, improves mathematical reasoning abilities while optimizing memory usage by replacing the Leave-One-Out method with a direct subtraction of the mean of all samples for a given prompt.
Offline preference optimization. In addition to DPO, several alternative preference optimization objectives have been proposed. IPO (Azar et al., 2023) addresses overfitting issues inherent in DPO. ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) aim to remove reliance on a reference model. R-DPO (Park et al., 2024) targets the reduction of exploitation due to sequence length, while KTO (Ethayarajh et al., 2024) handles preference optimization in the absence of pairwise data. CPO (Xu et al., 2024) and $\beta$-DPO (Wu et al., 2024b) focus on improving the quality of preference data. Another research direction addresses noise in offline alignment, which arises from the need to construct pairwise data. rDPO (Chowdhury et al., 2024), a variant of DPO, mitigates preference noise and enhances policy robustness, while DrDPO (Wu et al., 2024a) applies distributed robust optimization to tackle this issue. Other works have approached the problem through divergence regularization (Wang et al., 2024a; Huang et al., 2024), selection of high-quality data (Khaki et al., 2024; Ye et al., 2024), or reweighting loss functions (Wang et al., 2024c; Muldrew et al., 2024; Yang et al., 2024).
Iterative Preference Optimization. Offline preference optimization methods, such as DPO, face a limitation due to the lack of an explicit reward model, which hinders their ability to sample preference pairs from the optimal policy. To address this, iterative preference optimization techniques have been proposed. These methods iteratively update the reference model using the most recent policy model or generate new preference pairs in each iteration (Dong et al., 2024; Kim et al., 2024; Rosset et al., 2024; Xiong et al., 2024; Yuan et al., 2024; Chen et al., 2024b; Wu et al., 2024c; Gao et al., 2024). For instance, SPIN (Chen et al., 2024b) employs a self-play framework to fine-tune the model in a supervised manner, while Yuan et al. (2024) annotate preferences throughout the iterative process. REBEL improves sample quality by regressing the relative reward. Additionally, (Calandriello et al., 2024) generates data using a mixture policy, similar to the Nash-MD algorithm (Rosset et al., 2024).
Appendix B Proofs
B.1 Proof of Lemma 3.1
See 3.1
Proof.
We formally establish the gradient equivalence through pointwise convergence analysis. Let $\mathcal{D}$ be the data distribution and $\theta$ denote model parameters. Recall the gradient expressions:
SimPO Gradient:
$\nabla_{\theta}\mathcal{L}_{\mathrm{SimPO}}=-\beta\mathbb{E}_{\mathcal{D}}\left[\sigma\big{(}\beta(-M_{\theta}+\gamma)\big{)}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]$ (18)
RePO Gradient:
$\nabla_{\theta}\mathcal{L}_{\mathrm{RePO}}=-\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}(M_{\theta}<\gamma)\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]$ (19)
where $\sigma(\cdot)$ is the sigmoid function. The equivalence hinges on the limiting behavior of the sigmoid weighting term $s_{\theta}=\sigma(\beta(-M_{\theta}+\gamma))$. We analyze three cases:
Case 1: $M_{\theta}<\gamma$
Here, $-M_{\theta}+\gamma>0$. As $\beta\to\infty$,
$\lim_{\beta\to\infty}\sigma\big{(}\beta(-M_{\theta}+\gamma)\big{)}=\lim_{z\to\infty}\sigma(z)=1=\mathbb{I}(M_{\theta}<\gamma).$
Case 2: $M_{\theta}>\gamma$
Here, $-M_{\theta}+\gamma<0$. As $\beta\to\infty$,
$\lim_{\beta\to\infty}\sigma\big{(}\beta(-M_{\theta}+\gamma)\big{)}=\lim_{z\to-\infty}\sigma(z)=0=\mathbb{I}(M_{\theta}<\gamma).$
Case 3: $M_{\theta}=\gamma$
This occurs on a measure-zero set under continuous distributions. The limit becomes:
$\lim_{\beta\to\infty}\sigma(0)=\frac{1}{2}\neq\mathbb{I}(M_{\theta}<\gamma),$
which is negligible in expectation.
Thus, $\lim_{\beta\to\infty}s_{\theta}=\mathbb{I}(M_{\theta}<\gamma)$ almost everywhere. Substituting this into the SimPO gradient:
$\lim_{\beta\to\infty}\nabla_{\theta}\mathcal{L}_{\mathrm{SimPO}}$ $=-\lim_{\beta\to\infty}\beta\mathbb{E}_{\mathcal{D}}\left[s_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]$ (20)
[…]
To resolve the $\beta$ scaling, observe that for $M_{\theta}\neq\gamma$:
$\lim_{\beta\to\infty}\beta s_{\theta}=\begin{cases}\lim_{\beta\to\infty}\beta\cdot 1=\infty&\text{if }M_{\theta}<\gamma\\ \lim_{\beta\to\infty}\beta\cdot 0=0&\text{if }M_{\theta}>\gamma\end{cases}$ (22)
The divergence when $M_{\theta}<\gamma$ is mitigated by adaptive optimizers like Adam, which normalize gradient magnitudes through momentum terms. Formally, let $g_{\theta}=\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}$. Under Adam’s update rule:
$\theta_{t+1}=\theta_{t}-\eta\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon},$
where $\hat{m}_{t}$ and $\hat{v}_{t}$ are bias-corrected momentum estimates. The infinite gradient magnitude is absorbed into $\hat{m}_{t}/\sqrt{\hat{v}_{t}}$, effectively reducing to a unit-scaled update. Hence, in normalized update space:
$\lim_{\beta\to\infty}\beta s_{\theta}\cdot g_{\theta}\propto\mathbb{I}(M_{\theta}<\gamma)\cdot g_{\theta}.$
Combining these results:
$\lim_{\beta\to\infty}\nabla_{\theta}\mathcal{L}_{\mathrm{SimPO}}=-\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}(M_{\theta}<\gamma)\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]=\nabla_{\theta}\mathcal{L}_{\mathrm{RePO}},$ (23)
which completes the proof. ∎
B.2 Proof of Theorem 4.2
See 4.2
Proof.
We demonstrate that $h(x)=\frac{1}{a}\mathrm{ReLU}(-x)$ satisfies the convex envelope definition through three sequential arguments.
1. Convexity and Underestimation: The ReLU function is convex as the pointwise maximum of affine functions (Rule 3). The composition $h(x)=\frac{1}{a}\max(-x,0)$ preserves convexity through affine transformation (Rule 2). For all $x\in D$:
  • When $x<0$: $h(x)=-\frac{x}{a}\leq 1=\mathcal{L}_{0\text{-}1}(x)$, since $x\geq-a\implies-\frac{x}{a}\leq 1$
  • When $x\geq 0$: $h(x)=0=\mathcal{L}_{0\text{-}1}(x)$
Thus $h(x)\leq\mathcal{L}_{0\text{-}1}(x)$ over $D$.
2. Maximality Among Convex Underestimators: Let $g(x)$ be any convex function satisfying $g(x)\leq\mathcal{L}_{0\text{-}1}(x)$. For $x\in[-a,0)$, convexity implies:
$g(x)\leq\frac{-x}{a}g(-a)+\left(1+\frac{x}{a}\right)g(0)\leq\frac{-x}{a}$
since $g(-a)\leq 1$ and $g(0)\leq 0$. For $x\geq 0$, $g(x)\leq 0$. Hence $g(x)\leq h(x)$ for all $x\in D$.
3. Epigraph Characterization: The epigraph $\mathrm{epi}(h)$ coincides with the convex hull of $\mathrm{epi}(\mathcal{L}_{0\text{-}1})\cap(D\times\mathbb{R})$. The affine segment $h(x)=-\frac{x}{a}$ on $[-a,0)$ connects the points $(-a,1)$ and $(0,0)$, forming the tightest convex fit to the 0-1 loss’s discontinuity. By Theorem 1 in (Princeton University, Accessed 2025), this construction achieves the convex envelope. ∎
B.3 Proof of Corollary 4.3
See 4.3
Proof.
Part 1: By Theorem 1 in the lecture notes (Page 5), for any function $f$ and convex set $S$:
$\min_{x\in S}f(x)=\min_{x\in S}\mathrm{conv}_{S}f(x).$
Let $S=D$ and $f=\mathcal{L}_{0\text{-}1}$. The equality of minima implies:
$\{x^{*}\in D\mid\mathcal{L}_{0\text{-}1}(x^{*})=\min\mathcal{L}_{0\text{-}1}\}\subseteq\{x^{*}\in D\mid\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x^{*})=\min\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}\}.$
To show reverse inclusion, suppose $x^{*}\in\arg\min\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}$. Since $\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x^{*})\leq\mathcal{L}_{0\text{-}1}(x^{*})$ and $\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}$ attains its minimum at the same points as $\mathcal{L}_{0\text{-}1}$, $x^{*}$ must also minimize $\mathcal{L}_{0\text{-}1}$.
Part 2: For $D=[-a,b]$, both $\mathcal{L}_{0\text{-}1}(x)$ and $\frac{1}{a}\mathrm{ReLU}(-x)$ attain their minimum value 0 on $[0,b]$. For $x\in[-a,0)$, $\frac{1}{a}\mathrm{ReLU}(-x)$ is strictly decreasing, achieving its minimum at $x=0$. Thus:
$\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x)=[0,b].$
B.4 Proof of Corollary 4.5
See 4.5
Proof.
We demonstrate violation of the convex envelope’s defining property. Consider $D=[-1,1]$:
1. Underestimation Failure: For $x>0$:
$-\log\sigma(x)=-\log\left(\frac{1}{1+e^{-x}}\right)=\log(1+e^{-x})>0=\mathcal{L}_{0\text{-}1}(x)$
Thus $f_{\mathrm{log-sigmoid}}\not\leq\mathcal{L}_{0\text{-}1}$ over $D$, violating the envelope requirement.
2. Non-Maximality: Even if scaled, the logistic loss’s curvature differs from the ReLU envelope. For $x\in(-1,0)$, $\frac{d^{2}}{dx^{2}}(-\log\sigma(x))=\sigma(x)(1-\sigma(x))>0$, making it strictly convex – incompatible with the affine structure of $\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}$.
Hence $f_{\mathrm{log-sigmoid}}$ cannot be the convex envelope. ∎
Appendix C Experiments
C.1 Implementation Details
Empirical observations indicate significant performance sensitivity to model parameter initialization and learning rate selection across compared methods. To establish rigorous comparison benchmarks, we conducted systematic hyperparameter searches adhering to the specifications in each method’s original publication. The complete search space configuration is documented in Table 3. Notably, substantial architecture updates to both Llama3-8B and Instruct-7B necessitated re-implementation of the SimPO method, as the original implementation became incompatible with the revised model interfaces.
Training Protocol All experiments employed standardized training configurations to ensure comparability:
  • Batch size: 128 (consistent across methods)
  • Learning rate: Searched in {3e-7, 5e-7, 8e-7, 1e-6}
  • Training duration: Single epoch with cosine annealing schedule
  • Warmup: 10% of total training steps
  • Optimizer: Adam (Kingma & Ba, 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
  • Sequence length: 2048 tokens (fixed for all inputs)
The learning rate schedule follows a triangular policy with amplitude decay, selected through cross-validation on held-out development sets. All implementations utilize full-precision floating-point arithmetic to prevent gradient quantization artifacts.
Table 3: Various preference optimization objectives and hyperparameter search range.
Method Objective Hyperparameter
DPO (Rafailov et al., 2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in[0.01,0.05,0.1]$
IPO (Azar et al., 2023) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
CPO (Xu et al., 2024) $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
KTO (Ethayarajh et al., 2024) $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$
$\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\text{KL}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)\right]$ $\beta\in[0.01,0.05,0.1]$
ORPO (Hong et al., 2024) $-\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$ $\lambda\in[0.1,0.5,1.0,2.0]$
$\text{where}\,\,p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$
R-DPO (Park et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
$\beta\in[0.01,0.05,0.1]$
SimPO (Meng et al., 2024) $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in[2.0,4.0,6.0,8.0]$
$\gamma\in[0.3,0.5,1.0,1.2,1.4,1.6]$
RePO $\text{ReLU}[-(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma)]$ $\gamma\in[0.2,0.4,0.6,0.8]$
Table 4: The hyperparameter values in RePO used for each training setting.
Setting $\gamma$ Learning rate
Mistral-Instruct 0.4 6e-7
Llama3-Instruct 0.6 1e-6
Llama3-Instruct-v0.2 0.6 1e-6
Gemma2-Instruct 0.4 8e-7
Hyperparameters in RePO. Table 4 summarizes the hyperparameters utilized for RePO across different experimental settings. Our methodology only involves one hyperparameter: $\gamma$. Based on empirical evidence, we recommend setting $\gamma$ to a default value of $0.5$, as this configuration has consistently demonstrated reliability.
Decoding Hyperparameters. The decoding hyperparameters employed in this study align with those used in SimPO [^3] https://github.com/princeton-nlp/SimPO/tree/main/eval. We express our gratitude to the SimPO team for their generosity in sharing their insights and configurations, which have been instrumental in our work.
Computation Environment. All training experiments described in this paper were conducted using 8×A100 GPUs. The experimental setup follows the guidelines provided in the alignment-handbook repository [^4] https://github.com/huggingface/alignment-handbook, ensuring reproducibility and consistency with established practices.
C.2 Downstream Task Evaluation
Table 5: Downstream task evaluation results of tasks on the huggingface open leaderboard.
MMLU (5) ARC (25) HellaSwag (10) TruthfulQA (0) Winograd (5) GSM8K (5) Average
Llama3-Instruct
SFT 67.06 61.01 78.57 51.66 74.35 68.69 66.89
RRHF 67.20 61.52 79.54 53.76 74.19 66.11 67.05
SLiC-HF 66.41 61.26 78.80 53.23 76.16 66.57 67.07
DPO 66.88 63.99 80.78 59.01 74.66 49.81 65.86
IPO 66.52 61.95 77.90 54.64 73.09 58.23 65.39
CPO 67.05 62.29 78.73 54.01 73.72 67.40 67.20
KTO 66.38 63.57 79.51 58.15 73.40 57.01 66.34
ORPO 66.41 61.01 79.38 54.37 75.77 64.59 66.92
R-DPO 66.74 64.33 80.97 60.32 74.82 43.90 65.18
SimPO 65.63 62.80 78.33 60.70 73.32 50.72 65.25
RePO 64.95 62.03 77.58 60.96 72.93 66.49 67.49
Llama3-Instruct v0.2
SFT 67.06 61.01 78.57 51.66 74.35 68.69 66.89
RRHF 66.60 63.74 80.98 59.40 76.32 58.68 67.62
SLiC-HF 66.91 61.77 79.17 56.36 76.40 68.23 68.14
DPO 65.57 65.87 79.66 63.08 74.51 73.01 70.28
IPO 66.06 64.85 81.02 57.29 76.72 76.12 70.34
CPO 65.67 62.12 79.63 56.34 77.98 75.28 69.50
KTO 65.99 62.88 79.02 54.66 74.66 76.42 68.94
ORPO 65.75 63.99 79.91 57.02 78.06 75.13 69.98
R-DPO 66.17 65.36 79.98 57.94 75.06 75.36 69.98
SimPO 65.18 67.15 78.04 64.92 73.88 71.34 70.08
RePO 65.00 68.09 80.50 64.38 76.16 69.37 70.58
To assess the impact of RePO on downstream task performance, we evaluate models trained with different preference optimization methods on a diverse set of tasks from the Huggingface Open Leaderboard (Beeching et al., 2023). The tasks include MMLU (Hendrycks et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2022), Winograd (Levesque et al., 2012), and GSM8K (Cobbe et al., 2021). We adhere to standard evaluation protocols and present the results for all models in Table 5.
Overall Performance.
On average, RePO shows competitive performance across tasks, achieving an overall score of 67.49 on the Llama3-Instruct model and 70.58 on the Llama3-Instruct v0.2 model. The performance is generally close to that of other preference optimization methods, but it is worth noting that in some cases, it slightly lags behind models like SimPO or DPO, particularly on tasks such as ARC, HellaSwag, and TruthfulQA. However, the results suggest that RePO maintains a balanced performance profile across the evaluated tasks.
General Knowledge and Reasoning. On MMLU, which tests general knowledge and reasoning, RePO shows a slight reduction in performance (64.95 for Llama3-Instruct and 65.00 for Llama3-Instruct v0.2) compared to models such as RRHF and SimPO. This minor decline is consistent with the trend observed for other preference optimization methods and indicates that RePO may preserve general knowledge to a similar extent while possibly focusing more on improving performance in other areas such as reading comprehension and reasoning.
Reading Comprehension and Commonsense Reasoning. For ARC and HellaSwag, tasks related to reading comprehension and commonsense reasoning, RePO outperforms the base SFT model and exhibits competitive performance relative to other preference optimization methods. The Llama3-Instruct v0.2 model with RePO achieves a score of 80.50 on HellaSwag, which is comparable to the best-performing methods. This result suggests that RePO effectively improves the model’s ability to handle contextual understanding and reasoning, likely due to its optimization strategy.
Truthfulness. On the TruthfulQA task, RePO consistently shows improvements over the base SFT model, with a score of 60.96 for Llama3-Instruct and 64.38 for Llama3-Instruct v0.2. This indicates that RePO helps the model generate more truthful and reliable responses, aligning with trends observed in other preference optimization methods. The improvement in this area is especially notable given the inherent difficulty of this task, which tests the model’s ability to avoid generating false information.
Math Performance. The GSM8K benchmark, which tests mathematical reasoning, shows a drop in performance for RePO relative to the base SFT model. Specifically, the Llama3-Instruct model with RePO achieves a score of 66.49, which is lower than other methods such as SimPO or R-DPO, which focus more on improving mathematical reasoning. This drop is consistent with the trend observed across various preference optimization methods and may suggest that RePO is less effective in retaining mathematical reasoning abilities. Further investigation into this issue could provide insights into potential strategies for addressing this gap.
Task-Specific Variability. Overall, RePO exhibits varied performance across tasks. While it performs well in certain areas, such as commonsense reasoning and truthfulness, it lags behind in others, particularly in general knowledge (MMLU) and mathematical reasoning (GSM8K). This variability is in line with the performance trends observed for other preference optimization methods, which often show task-dependent improvements and declines. This suggests that RePO has strengths in some domains, but it may benefit from further refinement to improve performance across all tasks.
Conclusion and Future Work. In summary, RePO shows promising results across a range of downstream tasks, with notable improvements in tasks related to commonsense reasoning and truthfulness. However, it faces challenges in preserving general knowledge and mathematical reasoning. Future work could focus on incorporating additional strategies to mitigate performance drops in these areas, possibly through the integration of task-specific objectives or enhanced fine-tuning techniques. Additionally, further empirical studies are needed to explore the impact of different preference optimization datasets and objectives on downstream performance.
C.3 RePO with varying $\gamma$
Figure 8: Impact of the hyperparameter $\gamma$ on LC Win Rate and Raw Win Rate for Llama3-8B-Instruct (left) and Gemma2-9B-Instruct (right).
Figure 8 illustrates the effect of the hyperparameter $\gamma$ on model performance across two evaluation metrics: LC Win Rate and Raw Win Rate. The analysis is conducted on two models, Llama3-8B-Instruct (left) and Gemma2-9B-Instruct (right). The LC Win Rate, shown in red (left y-axis), represents the model’s alignment with learned preferences, whereas the Raw Win Rate, shown in blue (right y-axis), evaluates overall ranking performance based on human preference comparisons.
Moderate $\gamma$ values lead to optimal performance. Moderate values of $\gamma$ (0.4–0.6) yield the best balance between preference alignment and generalization. Both models achieve their highest LC Win Rate in this range, indicating that preference optimization is most effective when applied at an intermediate level. As $\gamma$ increases beyond 0.6, LC Win Rate starts to decline, likely due to overfitting, where the model overly aligns with preference data at the expense of generalization. Conversely, at $\gamma=0.0$, where no preference optimization is applied, the LC Win Rate remains low, emphasizing the necessity of preference tuning.
Raw Win Rate trends reveal model robustness differences. The Raw Win Rate follows a similar trend but highlights differences in robustness across models. For Llama3-8B-Instruct, the Raw Win Rate peaks at $\gamma=0.4$ before declining, suggesting that excessive preference optimization ($\gamma>0.6$) negatively impacts the model’s ability to generalize. In contrast, Gemma2-9B-Instruct exhibits a more stable Raw Win Rate across a wider range of $\gamma$, reaching its highest performance at $\gamma=0.6$ before experiencing a sharp decline at $\gamma=0.8$. This suggests that Gemma2-9B-Instruct maintains better robustness to preference optimization compared to Llama3-8B-Instruct.
Gemma2-9B-Instruct outperforms Llama3-8B-Instruct. Gemma2-9B-Instruct consistently outperforms Llama3-8B-Instruct in both LC Win Rate and Raw Win Rate. This observation indicates that Gemma2-9B-Instruct not only aligns more effectively with learned preferences but also retains superior generalization capability. The results highlight the importance of carefully selecting $\gamma$ to avoid performance degradation at extreme values. Future work could explore adaptive strategies for dynamically tuning $\gamma$, ensuring that preference optimization enhances alignment without compromising generalization.
C.4 Analsis on RePO++
Figure 9 illustrates the relationship between the gradient and the implicit reward margin. As shown in the figure, when the implicit reward margin is greater than $\gamma$, the gradient becomes zero. In this case, the model can stop updating for well-separated pairs, thus preventing overfitting. On the other hand, when the implicit reward margin is less than $\gamma$, the model continues to increase the weight for less-separated pairs. Furthermore, the harder the pair is to distinguish, the larger the gradient becomes, eventually converging to 1.0. This behavior is reminiscent of curriculum learning, where more difficult samples are assigned higher weights.
Figure 9: Gradient weighting functions of RePO++ ($s_{\theta}\cdot\mathbb{I}(M_{\theta}<\gamma)$).

### v2

@@
-RePO: ReLU-based Preference Optimization
+RePO: Understanding Preference Learning Through ReLU-Based Optimization
 Abstract
-Aligning large language models (LLMs) with human preferences is critical for real-world deployment, yet existing methods like RLHF face computational and stability challenges.
-While DPO establishes an offline paradigm with single hyperparameter $\beta$, subsequent methods like SimPO reintroduce complexity through dual parameters ($\beta$, $\gamma$).
-We propose ReLU-based Preference Optimization (RePO), a streamlined algorithm that eliminates $\beta$ via two advances: (1) retaining SimPO’s reference-free margins but removing $\beta$ through gradient analysis, and (2) adopting a ReLU-based max-margin loss that naturally filters trivial pairs.
-Theoretically, RePO is characterized as SimPO’s limiting case ($\beta\to\infty$), where the logistic weighting collapses to binary thresholding, forming a convex envelope of the 0-1 loss.
-Empirical results on AlpacaEval 2 and Arena-Hard show that RePO outperforms DPO and SimPO across multiple base models, requiring only one hyperparameter to tune.
-LLM Alignment, RLHF, Data Selection
+Preference learning has become a common approach in various recent methods for aligning large language models with human values.
+These methods optimize the preference margin between chosen and rejected responses, subject to certain constraints for avoiding over-optimization.
+In this paper, we report surprising empirical findings that simple ReLU activation can learn meaningful alignments even using none of the following: (i) sigmoid-based gradient constraints, (ii) explicit regularization terms.
+Our experiments show that over-optimization does exist, but a threshold parameter $\gamma$ plays an essential role in preventing it by dynamically filtering training examples.
+We further provide theoretical analysis demonstrating that ReLU-based Preference Optimization (RePO) corresponds to the convex envelope of the 0-1 loss, establishing its fundamental soundness.
+Our “RePO” method achieves competitive or superior results compared to established preference optimization approaches.
+We hope this simple baseline will motivate researchers to rethink the fundamental mechanisms behind preference optimization for language model alignment.
 1 Introduction
-Aligning large language models (LLMs) with human preferences is essential for their effective deployment in real-world applications (Anil et al., 2023; Touvron et al., 2023; OpenAI, 2023; Bubeck et al., 2023), ensuring that their outputs are beneficial, accurate, and adhere to human values (Ouyang et al., 2022) while minimizing potential safety risks.
-A primary approach for achieving this alignment is Reinforcement Learning from Human Feedback (RLHF).
-It typically entails a multi-stage process: initially, a reward model is trained using ranked preference data; subsequently, the LLM is optimized with the guidance of reward model, using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017).
-Despite its effectiveness, RLHF faces significant challenges, such as high computational costs and training instability (Rafailov et al., 2023; Zhao et al., 2023).
-Recent research has been exploring simpler offline alternatives.
-Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one such approach.
-It reparameterizes the reward function in RLHF, bypassing the explicit learning of the reward model, and thus enabling direct training of LLMs with human preference data.
-Formally, the DPO objective is a logistic-log loss, which applies a log-sigmoid activation to the reward margin between the preferred and less-preferred responses to the same input.
-Here, the reward is derived from comparing the policy model of the LLM with a reference model.
-Building on this, follow-up studies, especially SLiC-HF (Zhao et al., 2023) and SimPO (Meng et al., 2024), focus on further simplifications on the loss function and reward margins separately.
-Specifically, SLiC-HF replaces the logistic-log loss in DPO with a hinge loss, achieving a faster-decaying tail.
-SimPO, on the other hand, introduces reference-free reward margins with logistic-log loss, offering a more streamlined and efficient approach.
-While advancing the field and achieving leading performance, they introduce additional complexity on hyperparameter tuning.
-SimPO inherits DPO’s reward margin strength controller $\beta$ and adds a target reward margin $\gamma$.
-Similarly, SLiC-HF utilizes $\gamma$, but also introduces a regularization weight $\lambda$.
-The necessity of dual-tuning these hyperparameters poses significant challenges in model training and deployment.
-This leads to a natural question:
-Can we develop a simpler offline preference optimization algorithm?
-Figure 1: RePO, SimPO, and DPO primarily differ in their loss functions, as highlighted in the shaded box.
-RePO either outperforms or achieves comparable performance to DPO and SimPO across various settings on AlpacaEval 2, with only a single hyperparameter $\gamma$.
-“IT” denotes pre-trained instruction-tuned models.
-e.g., “Gemma2-IT 9B” refers to Gemma2-9B (instruct setup).
-In this paper, we propose a simple idea: simultaneously simplifying the log-sigmoid activation and reward margins, to achieve comparable or even better performance.
-For reward margins, we adopt the reference-free approach of SimPO but eliminate the hyperparameter $\beta$.
-For the log-sigmoid activation, we employ ReLU to the reference-free reward margins, following a max-margin (support vector machine) paradigm (Boser et al., 1992; Cortes & Vapnik, 1995), similar to SLiC-HF.
-Consequently, our method can be interpreted as “SimPO without $\beta$” and “SLiC-HF without $\lambda$”.
-Interestingly, it bridges these methods by omitting a core component from each while maintaining competitive performance, demonstrating the effectiveness of this streamlined approach.
-We term this ReLU-based Preference Optimization RePO.
-Theoretically, we demonstrate that RePO emerges as the natural endpoint of SimPO’s architectural evolution: when $\beta\to\infty$, SimPO’s logistic weighting collapses to RePO’s binary thresholding (Lemma 3.1).
-This asymptotic behavior motivates our key theoretical insight — the ReLU loss in RePO is precisely the convex envelope of the 0-1 loss (Theorem 4.2), inheriting its global optimality while enabling tractable gradient-based optimization.
-Practically, this formulation provides the following advantages:
-  • Simplified Hyperparameter Tuning: RePO removes the hyperparameter $\beta$, requiring only the tuning of $\gamma$ within the range $[0,1]$.
-Empirical results show that RePO can achieve performance comparable to SimPO on three models (Llama, Mistral, and Gemma) even with a fixed $\gamma=0.5$, significantly reducing the complexity of hyperparameter tuning.
-  • Effective Data Filtering: By using ReLU instead of a continuous sigmoid function, RePO interprets the target reward margin $\gamma$ as a strict “cutoff point”.
-This mechanism zeroes out gradients for preference pairs whose reward margin exceeds $\gamma$ (well-separated pairs), while assigning unit weight to less-separated pairs.
-This selective approach minimizes overfitting by focusing optimization on challenging data points.
-  • Controllable Over-Optimization: The “cutoff point” $\gamma$ controls the mean of reward margin within a batch, representing a new metric for evaluating over-optimization.
-Experimental results suggest that this metric correlates with model behavior and can replace KL divergence as a simpler alternative, potentially controlling optimization without requiring a reference model.
-While RePO effectively filters out well-separated pairs by zeroing their gradients, it assigns a unit gradient weight to all less-separated pairs.
-This implicit uniform weighting overlooks the varying importance of these pairs, potentially limiting its ability to prioritize more informative examples.
-To address this limitation, we extend the concept of RePO within SimPO and propose a new method, RePO++, by leveraging SimPO’s logistic-log loss to differentiate the importance of less-separated pairs.
-This approach assigns greater weights to less-separated pairs, while continuing to filter out well-separated pairs.
-Notably, only a single line of code is required to seamlessly integrate RePO into SimPO, resulting in RePO++.
-In a nutshell, our main contribution is ReLU-based Preference Optimization (RePO), a simple but effective reference-free algorithm for training language models from preferences.
-With its straightforward motivation, RePO is grounded in a principled theoretical framework and consistently outperforms DPO and SimPO across three base model settings (Mistral2-7B, Llama3-8B, and Gemma2-9B) on AlpacaEval 2 (cf.
+Recent years have witnessed significant advances in aligning large language models (LLMs) with human preferences [1, 2, 3, 4].
+A primary approach, Reinforcement Learning from Human Feedback (RLHF) [5], first trains a reward model on preference data and then optimizes the LLM via reinforcement learning.
+While effective, RLHF’s computational costs and training instability [6, 7] have motivated simpler offline alternatives like Direct Preference Optimization (DPO) [6], which bypasses explicit reward modeling.
+Take DPO as a representative example: it optimizes the alignment margin between a preferred and a less-preferred response to the same prompt, as Figure 1 shows.
+The alignment of each response is quantified via an implicit reward, defined as the log-ratio of the predicted likelihoods under the policy model (i.e., the LLM being optimized) and a reference model (e.g., a fixed supervised fine-tuned (SFT) model).
+A fundamental challenge in preference learning is over-optimization — where models excessively amplify reward margins between preferred and non-preferred responses, potentially degrading generation quality [8, 9, 10].
+Several approaches have been developed to mitigate this issue.
+DPO [6] and SimPO [11] employ sigmoid weighting through log-sigmoid activation that diminishes gradients as reward margins increase, naturally preventing over-optimization.
+The $\beta$ parameter controls gradient distribution sharpness — larger values produce more binary-like gradients, as illustrated in Figure 2.
+SLiC-HF [7] addresses over-optimization differently by incorporating an SFT regularization term that anchors the model to its initial policy [12], preventing excessive drift toward maximizing preference signals.
+These mechanisms effectively balance preference optimization with generation quality preservation, forming the foundation of current preference learning approaches.
+Here, we present a surprising empirical finding: a simple ReLU activation can work well with none of the above strategies for mitigating over-optimization.
+Our analysis reveals that as parameter $\beta$ in SimPO approaches infinity, its sigmoid weighting naturally converges to a binary thresholding mechanism — motivating our exploration of ReLU-based Preference Optimization (RePO).
+This mechanism uses a single ReLU function with only one hyperparameter $\gamma$, creating a clear decision boundary that selectively updates sample pairs with insufficient reward margins ($M_{\theta}<\gamma$) while filtering out well-separated pairs ($M_{\theta}\geq\gamma$).
+We illustrate this “RePO” method in Figure 1.
+Figure 1: Comparing preference learning mechanisms.
+RePO employs a simpler binary thresholding mechanism than SimPO and DPO, as highlighted in the shaded box.
+Despite its simplicity, this mechanism achieves competitive results by naturally preventing over-optimization.
+Figure 2: Gradient weighting functions of SimPO ($s_{\theta}$) and RePO ($\mathbb{I}(M_{\theta}<\gamma)$).
+As $\beta\to\infty$, $s_{\theta}$ converges to the binary indicator (red line), establishing RePO as the limit case of SimPO.
+Thanks to its conceptual simplicity, RePO can serve as a hub that relates several existing methods.
+In essence, our method can be viewed as “SimPO without log-sigmoid” or “SLiC-HF without SFT regularization term”.
+Interestingly, RePO is related to each method by removing one of its core components.
+Even so, RePO effectively prevents over-optimization while performing competitively or better (cf.
 Figure 1).
-We hope our exploration will motivate people to rethink the fundamental roles of loss function design for preference optimization.
+We empirically show that over-optimization do exist, but ReLU activation with threshold $\gamma$ is critical to prevent such solutions.
+This implies that in over-optimization regimes, selecting which examples to learn from is more critical than determining how much to learn from each.
+The $\gamma$ threshold induces an emergent data filtering behavior, focusing dynamically on challenging samples relative to the model’s current capability.
+Our theoretical analysis reveals that RePO’s ReLU loss corresponds precisely to the convex envelope of the 0-1 loss (Theorem 4.2), explaining why such a simple mechanism is so effective.
+Our simple baseline suggests that the ReLU activation with a proper threshold $\gamma$ can be an essential reason for the common success of related methods.
+We believe this work’s significance lies in revealing how preference learning principles may be simpler than previously thought.
+By questioning conventional wisdom about necessary components, we hope to motivate researchers to reconsider the fundamental mechanisms behind preference optimization.
 2 Preliminaries
-Offline Alignment.
-In the offline alignment setting, we utilize a dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$, where each entry consists of a prompt $x$ accompanied by two responses: $y_{w}$ is the preferred (winning) response, and $y_{l}$ is the less preferred (losing) response, both generated by a reference policy $\pi_{\text{ref}}$.
-The primary challenge is to find a latent reward function $r^{*}(x,y)$ that governs these preference relationships.
-Although $r^{*}(x,y)$ is not directly observable, the Bradley-Terry (BT) model (Bradley & Terry, 1952) provides a probabilistic framework for modeling the probability of $y_{w}$ over $y_{l}$ given $x$:
-$\mathbb{P}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r^{*}(x,y_{w}))}{\exp(r^{*}(x,y_{w}))+\exp(r^{*}(x,y_{l}))}.$ (1)
-Here, $r^{*}(x,y)$ represents the latent reward for a response $y$ in the context of prompt $x$.
-The aim of offline alignment is to develop a policy $\pi_{\theta}$ capable of accurately estimating this latent reward function using the dataset $\mathcal{D}$.
-Reinforcement Learning from Human Feedback (RLHF).
-Traditional offline alignment methods often leverage reinforcement learning augmented with a KL-divergence regularization term in the reward objective (Bai et al., 2022; Ziegler et al., 2019; Ouyang et al., 2022).
-The optimization objective, parameterized by a regularization parameter $\beta>0$, is expressed as:
-$\mathop{\max}_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y\mid x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)),$ (2)
-where $r_{\phi}(x,y)$ represents the reward function trained using the BT model on the preference data, $\pi_{\theta}$ is the policy being optimized, and $\pi_{\text{ref}}$ is the fixed reference policy, typically derived from supervised fine-tuning.
-The KL-divergence term ensures that $\pi_{\theta}$ remains close to $\pi_{\text{ref}}$, mitigating risks of excessive divergence during training.
 Directed Preference Optimization (DPO).
-DPO (Rafailov et al., 2023) stands out as a leading method for offline preference optimization by eliminating the need for an explicit reward model.
-Instead, it reformulates the reward function $r(x,y)$ as a closed-form expression based on policy ratios:
+DPO [6] stands out as a leading method for offline preference optimization by eliminating the need for an explicit reward model.
+Instead, it reformulates the reward $r(x,y)$ as a closed-form expression based on policy ratios:
 $r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x),$ (1)
 where $Z(x)$ is a partition function that does not depend on $y$. This leads to the DPO loss for a given triplet $(x,y_{w},y_{l})$ as:
-$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}$ (4)
-$\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right)\right],$
+$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right)\right],$ (2)
 where $\sigma(\cdot)$ denotes the sigmoid function. This loss encourages the policy $\pi_{\theta}$ to prefer $y_{w}$ over $y_{l}$ in alignment with the reference policy.
 Sequence Likelihood Calibration (SLiC-HF).
-SLiC-HF (Zhao et al., 2023) advances preference optimization with two key innovations: (1) it employs a sequence-level calibration loss that contrasts the log-probability difference between preferred and dispreferred responses using a margin $\gamma$, and (2) it integrates a regularization term to prevent divergence from the SFT policy, avoiding the need for an explicit KL penalty.
+SLiC-HF [7] advances preference optimization with two key innovations: (1) it employs a sequence-level calibration loss that contrasts the log-probability difference between preferred and dispreferred responses using a margin $\gamma$, and (2) it integrates a regularization term to prevent divergence from the SFT policy, avoiding the need for an explicit KL penalty.
 The SLiC-HF loss function is defined as:
-$\mathcal{L}_{\text{SLiC-HF}}(\pi_{\theta})=$ $\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\Big{[}\text{ReLU}\Big{(}-\Big{(}\log{\pi_{\theta}(y_{w}\mid x)}$ (5)
-$-\log\pi_{\theta}(y_{l}\mid x)-\gamma\Big{)}\Big{)}-\lambda\log{\pi_{\theta}(y_{w}\mid x)}\Big{]}.$
+$\mathcal{L}_{\text{SLiC-HF}}(\pi_{\theta})=$ $\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\Big[\text{ReLU}\Big(-\Big(\log{\pi_{\theta}(y_{w}\mid x)}-\log\pi_{\theta}(y_{l}\mid x)-\gamma\Big)\Big)-\lambda\log{\pi_{\theta}(y_{w}\mid x)}\Big].$ (3)
 Simple Preference Optimization (SimPO).
-SimPO (Meng et al., 2024) advances preference optimization with two key innovations: (1) it normalizes the reward by the length of the response, calculating the average log-probability per token for a response under the policy $\pi_{\theta}$, and (2) it incorporates a target reward margin $\gamma$ to ensure that the reward difference between the preferred and less preferred responses exceeds this margin.
+SimPO [11] advances preference optimization with two key innovations: (1) it normalizes the reward by the length of the response, calculating the average log-probability per token for a response under the policy $\pi_{\theta}$, and (2) it incorporates a target reward margin $\gamma$ to ensure that the reward difference between the preferred and less preferred responses exceeds this margin.
 The SimPO loss function is defined as:
-$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}$ (6)
-$\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|}-\gamma\right)\right)\right],$
+$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|}-\gamma\right)\right)\right],$ (4)
 where $|y|$ denotes the number of tokens in response $y$, ensuring length-aware scaling of rewards, and $\gamma$ is the predefined margin that enforces a minimum difference in rewards between $y_{w}$ and $y_{l}$.
 To align with subsequent discussions, we modify the original SimPO formulation by setting $\gamma$ to $\gamma/\beta$.
-3 Analysis on RePO
-In this section, we first present the expression for RePO and compare its performance with that of SimPO through experiments.
-We then analyze the advantages and key factors behind the success of RePO from the perspectives of gradient behavior, data filtering, and over-optimization.
-3.1 RePO
-Method Formulation.
-We propose RePO by simplifying the SimPO loss function through two key modifications: (i) removing the hyperparameter $\beta$, and (ii) replacing the log-sigmoid function with a ReLU activation.
-First, we define the implicit reward margin $M_{\theta}$ as:
+3 Exploring Simple ReLU Activation in Preference Learning
+In this section, we explore what makes a simple ReLU activation function effective for preference learning.
+We first examine the surprising relationship between ReLU activation and sigmoid weighting through empirical experiments.
+Then, we investigate the key properties that emerge from this simple mechanism, specifically through the lens of gradient behavior, data filtering patterns, and over-optimization control.
+3.1 Examining ReLU-based Preference Optimization
+Simplification exploration.
+Our exploration began by questioning whether log-sigmoid activation or SFT regularization are truly necessary for mitigating over-optimization.
+We simplified the SimPO loss function through two key modifications: (i) removing the hyperparameter $\beta$, and (ii) replacing the log-sigmoid function with a ReLU activation.
+We adopt the length-normalized implicit reward margin $M_{\theta}$ (as introduced in SimPO [11]):
 $M_{\theta}=\frac{\log\pi_{\theta}(y_{w}\mid x)}{|y_{w}|}-\frac{\log\pi_{\theta}(y_{l}\mid x)}{|y_{l}|},$ (5)
-which measures the policy’s relative preference between responses. Using $M_{\theta}$, the proposed loss function is:
+which quantifies the policy’s preference between responses. Using $M_{\theta}$, we examine a loss function with the following form:
 $\mathcal{L}_{\text{RePO}}(\pi_{\theta})=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\left[\mathrm{ReLU}\left(-(M_{\theta}-\gamma)\right)\right],$ (6)
 where $\gamma\in[0,1]$ is the sole hyperparameter representing the target reward margin.
-Comparison between RePO and SimPO / SLiC-HF.
-As formalized in Equation 8, RePO simplifies the SimPO framework by replacing the log-sigmoid activation with a ReLU operator and removing the need to tune the hyperparameter $\beta$.
-Simultaneously, it eliminates the explicit regularization term in SLiC-HF, thereby removing the associated hyperparameter $\lambda$.
-This dual simplification makes RePO a unified specialization of both SimPO and SLiC-HF, reducing hyperparameter tuning complexity while maintaining competitive performance, as demonstrated in Section 3.2.
-Training Dynamics Monitoring. We track optimization progress with two metrics:
-  • Batch-level reward margin:
-    $m_{\text{batch}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{B}}[M_{\theta}],$ (9)
-    which reflects the average reward margin within each training batch $\mathcal{B}$.
-  • Dataset-level reward margin:
-    $m_{\mathcal{D}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}[M_{\theta}],$ (10)
-    which measures the overall reward margin on the entire training set $\mathcal{D}$.
-Here, $m_{\text{batch}}$ provides immediate feedback during each update, while $m_{\mathcal{D}}$ captures the global preference strength of the policy over the complete dataset.
-Gradient Analysis. We analyze the gradient dynamics of RePO and SimPO to highlight their fundamental differences, the gradients are as follows:
+Gradient behavior investigation. We examine the gradient dynamics of RePO and SimPO to reveal how our simplified approach addresses over-optimization:
 $\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\beta\mathbb{E}_{\mathcal{D}}\left[s_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right],$ (7)
@@
 The terms $\nabla_{\theta,y_{w}}=\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}\mid x)$ and $\nabla_{\theta,y_{l}}=\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}\mid x)$ correspond to the gradients that increase the probability of the “winning” response $y_{w}$ and decrease the probability of the “losing” response $y_{l}$, respectively.
-The scaling factor $\beta$ in Equation 11 linearly amplifies gradient magnitudes but does not alter the relative update directions in adaptive optimizers like Adam (Kingma & Ba, 2014), as the momentum terms automatically normalize scale variations.
+The scaling factor $\beta$ in Equation 7 linearly amplifies gradient magnitudes but does not alter the relative update directions in adaptive optimizers like Adam [13], as the momentum terms automatically normalize scale variations.
 We therefore omit $\beta$ in Figure 2 for clearer visualization of the weighting function shapes.
-RePO’s gradient (Equation 12) applies uniform updates to pairs with $M_{\theta}<\gamma$, while SimPO (Equation 11) employs $\beta$-scaled continuous weighting through $s_{\theta}$.
-Figure 2 visualizes this contrast, showing how RePO’s binary thresholding emerges as a limiting case of SimPO when $\beta\to\infty$.
-Figure 2: Gradient weighting functions of SimPO ($s_{\theta}$) and RePO ($\mathbb{I}(M_{\theta}<\gamma)$).
-As $\beta\to\infty$, $s_{\theta}$ converges to the binary indicator (dashed line), establishing RePO as the limit case of SimPO.
+The key insight is that RePO’s ReLU-based gradient (Equation 8) applies uniform updates only to pairs with $M_{\theta}<\gamma$, while SimPO’s gradient (Equation 7) uses continuous $\beta$-scaled weights.
+Figure 2 visualizes this difference, showing RePO as the limiting case of SimPO as $\beta\to\infty$.
 Lemma 3.1 (Gradient Equivalence in the SimPO-to-RePO Limit).
@@
 Please check Appendix for all proofs.
-Lemma 3.1 establishes RePO as the asymptotic limit of SimPO with large $\beta$.
-This explains why RePO achieves comparable performance to SimPO (Section 3.2) while eliminating $\beta$ tuning.
-The binary thresholding in RePO also induces implicit data filtering, which we analyze in Section 3.3.
-3.2 Relationship between SimPO and RePO
+Lemma 3.1 establishes RePO as the asymptotic limit of SimPO with large $\beta$, explaining two key advantages we will demonstrate in Section 3.2: comparable performance without $\beta$ tuning complexity, and an effective binary thresholding mechanism that induces implicit data filtering for controlling over-optimization.
+3.2 Empirical Study
 The previous section analyzes the relationship between SimPO and RePO from the perspective of gradient behavior. In this section, we compare their performance from an empirical standpoint.
 Experimental setup.
-We adopt the experimental setup used in SimPO (Meng et al., 2024), with Llama3-8B and Gemma2-9B as our base models.
-For comparison, we use pre-trained instruction-tuned models as reference policies.
-To ensure consistency, we employ the same training datasets as in SimPO: princeton-nlp/llama3-ultrafeedback-armorm [^1] https://huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm for Llama3-8B and princeton-nlp/gemma2-ultrafeedback-armorm [^2] https://huggingface.co/datasets/princeton-nlp/gemma2-ultrafeedback-armorm for Gemma2-9B.
-In all experiments involving SimPO, we set $\beta=10.0$ and $\gamma=0.4$ for Gemma2-9B and $\beta=10.0$ and $\gamma=0.3$ for Llama3-8B, unless stated otherwise.
+We evaluate this approach using SimPO’s experimental setup [11] with Llama3-8B and Gemma2-9B models (Instruct setup).
+For consistency, we use the same training datasets as SimPO: princeton-nlp/llama3-ultrafeedback-armorm for Llama3-8B and princeton-nlp/gemma2-ultrafeedback-armorm for Gemma2-9B.
+For all SimPO experiments, we set $\beta=10.0$ and $\gamma=0.4$ for Gemma2-9B and $\beta=10.0$ and $\gamma=0.3$ for Llama3-8B, unless otherwise specified.
+We track optimization progress using two reward margin metrics:
+$m_{\text{batch}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{B}}[M_{\theta}],\quad m_{\mathcal{D}}=\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}[M_{\theta}],$ (10)
+measuring response separation within each batch ($m_{\text{batch}}$) and across the entire training set ($m_{\mathcal{D}}$).
 Figure 3: Performance of SimPO with varying $\beta$ and RePO on AlpacaEval2 benchmark.
 Evaluation benchmarks.
-We evaluate our models on two widely used open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024).
-These benchmarks assess the models’ conversational abilities across various queries.
-For AlpacaEval 2, we report both the length-controlled win rate (LC) and raw win rate (WR).
-For Arena-Hard, we report the win rate (WR).
-Observation 1: Large $\beta$ enhances SimPO’s performance.
-We first experiment with SimPO by varying the parameter $\beta$ over the set {1.0, 3.0, 5.0, 10.0}, while keeping $\gamma$ fixed.
-As shown in Figure 3, increasing $\beta$ consistently improves SimPO’s performance,with convergence observed as $\beta$ grows larger
-Observation 2: RePO achieves competitive performance.
-RePO achieves performance comparable to SimPO with a large $\beta$.
-As shown in Figure 3, RePO achieves win rates of 51.1% on Llama3-8B and 66.6% on Gemma2-9B, comparable to SimPO’s performance.
-This aligns with Lemma 3.1, which establishes that RePO can be interpreted as a limiting case of SimPO as $\beta\to\infty$.
-Figure 4: Line plot illustrating the variation in RePO’s performance with different values of $\gamma$, measured by the AlpacaEval 2 LC Win Rate.
-The accompanying bar chart shows the mean implicit reward margin, $m_{\mathcal{D}}$, computed across all training pairs.
-Observation 3: $\gamma$ is a crucial hyperparameter.
-In RePO, the gradients become zero when the implicit reward margin exceeds $\gamma$, indicating that the model $\pi_{\theta}$ is confident in its predictions for these data pairs.
-These pairs, considered well-separated, are excluded from updates, mitigating overfitting.
-Thus, $\gamma$ serves as a critical “cutoff point”.
-As shown in Figure 4, the performance of RePO is highly sensitive to $\gamma$.
-Increasing $\gamma$ initially improves performance, but beyond a certain threshold, performance degrades.
-Intuitively, A small $\gamma$ retains many zero-gradient data points, leading to under-filtering.
-Conversely, a large $\gamma$ causes many data points to have unit gradient weight, potentially leading to overfitting.
-Further discussion on $\gamma$’s role in data filtering is provided in subsequent sections.
+We evaluate on two established benchmarks for open-ended generation: AlpacaEval 2 [14] (measuring instruction-following quality against GPT-4) and Arena-Hard [15] (testing complex reasoning).
+For AlpacaEval 2, we report both length-controlled win rate (LC-Win Rate) and raw win rate (WR); for Arena-Hard, we report the standard win rate.
+Observation 1: Large $\beta$ enhances SimPO’s performance when paired with appropriate $\gamma$.
+We systematically evaluate SimPO across varying values of $\beta\in\{1.0,3.0,5.0,10.0\}$, while maintaining fixed $\gamma$ values that we empirically determined to be suitable for each model architecture ($\gamma=0.4$ for Gemma2-9B and $\gamma=0.3$ for Llama3-8B).
+As shown in Figure 3, increasing $\beta$ leads to consistent performance improvements across all evaluation metrics, with diminishing returns observed beyond $\beta=5.0$.
+These findings align with observations in the SimPO paper [^1] In their official repository, the authors note: “SimPO requires a much larger $\beta$ than DPO… In many cases, an even larger (e.g., 10) could yield better results.”.
 Figure 4: Implicit reward margin $M_{\theta}$ distribution across training steps (total: 467) for RePO at $\gamma=0.4$.
@@
 Numbers: fraction of samples above $\gamma$.
-3.3 Data Filtering by RePO
-Building upon the preceding discussion, we observe that adjusting the parameter $\gamma$ defines a “cutoff point” that determines which sample gradients are set to zero, effectively implementing data filtering.
-In this section, we analyze the training process of RePO through the lens of data filtering.
-Observation 4: $\gamma$ controls the distribution of implicit reward margins.
-We also report $m_{\mathcal{D}}$ (cf.
-Equation 10), the mean of implicit reward margins across all training pairs.
-As shown in the bar chart of Figure 4, as $\gamma$ increases, $m_{\mathcal{D}}$ rises.
-However, alignment performance w.r.t.
-LC improves initially but declines once $\gamma$ exceeds a certain threshold.
-This suggests that while a larger $\gamma$ increases the separation between training pairs, excessive separation can introduce the risk of reward hacking (Amodei et al., 2016), leading to a performance drop.
-Observation 5: Less data, better alignment.
-Figure 5 illustrates the distribution of the implicit reward margin $M_{\theta}$ across different training stages, alongside the ratio of zero gradients (i.e., when $M_{\theta}$ exceeds $\gamma$).
-As training progresses, the model’s ability to discriminate between winning and losing samples improves, resulting in a steady increase in both the implicit reward margin and the ratio of filtered data.
-Notably, the filtered data ratio rises from 48% to 58% between steps 200 and 400.
-Despite using only half of the samples for gradient updates, the model achieves optimal performance, indicating that RePO effectively utilizes a smaller, more informative subset of training data.
+Figure 5: Line plot of RePO performance (AlpacaEval 2 LC Win Rate) and bar chart of mean reward margins ($m_{\mathcal{D}}$) across varying $\gamma$ values. See Appendix D.3 for details.
+Observation 2: RePO matches high-$\beta$ SimPO.
+RePO achieves performance comparable to SimPO with a large $\beta$.
+As shown in Figure 3, RePO achieves win rates of 51.1% on Llama3-8B and 66.6% on Gemma2-9B, comparable to SimPO’s performance.
+This aligns with Lemma 3.1, which establishes that RePO can be interpreted as a limiting case of SimPO as $\beta\to\infty$.
+Observation 3: $\gamma$ threshold creates a natural alignment-optimization tradeoff.
+Our experiments track the mean implicit reward margin $m_{\mathcal{D}}$ (cf.
+Equation 10) across training pairs.
+As Figure 5 illustrates, increasing $\gamma$ directly elevates $m_{\mathcal{D}}$ while performance follows an inverted U-shaped pattern — improving initially but declining beyond a critical threshold.
+In RePO, gradients vanish when the implicit reward margin exceeds $\gamma$, effectively filtering out well-separated pairs from updates.
+This mechanism creates a fundamental tradeoff: small $\gamma$ values retain excessive zero-gradient samples causing under-filtering, while large $\gamma$ values force updates on most samples, potentially leading to over-optimization [8] and ultimately degrading performance.
+Observation 4: RePO creates a natural learning curriculum via progressive filtering.
+Figure 4 reveals an unexpected pattern in how the distribution of implicit reward margins $M_{\theta}$ evolves throughout training.
+As learning progresses, the model’s ability to discriminate between winning and losing samples naturally improves, resulting in a steady increase in both the implicit reward margin and the ratio of filtered data.
+Notably, the filtered data ratio rises from 13% to 58% between steps 100 and 400.
+This creates an emergent curriculum where the model initially learns from a broader set of examples and gradually focuses on the more challenging ones — despite using only half of the samples for gradient updates in later stages, the model achieves optimal performance.
 3.3 Over-Optimization Analysis
-When a larger $\gamma$ is selected, RePO permits most data to bypass the filter, resulting in a continuous increase in the reward margin between winning and losing samples.
-This phenomenon, a manifestation of reward exploitation or hacking (Amodei et al., 2016), is often referred to as “over-optimization” in the RLHF literature, and has been empirically investigated in both controlled experiments (Gao et al., 2023) and user studies (Dubois et al., 2023).
+The study of over-optimization can be traced back to traditional RLHF literature, and has been empirically investigated in both controlled experiments [9] and user studies [10].
+In this work, we follow their experimental setup to further explore this phenomenon.
 Model Over-Optimization: Building on Rafailov et al.
-(2024), we investigate over-optimization in RePO, by evaluating six different values of $\gamma$ ($0.0,0.2,0.4,0.6,0.8,1.0$), each corresponding to varying levels of data filtering.
+[16], we investigate over-optimization in RePO, by evaluating six different values of $\gamma$ ($0.0,0.2,0.4,0.6,0.8,1.0$), each corresponding to varying levels of data filtering.
 Across all cases, we observe a distinct hump-shaped performance pattern: while moderate filtering improves alignment, excessive filtering causes performance to degrade, highlighting the over-optimization effect.
-Figure 6: Over-optimization results for RePO on Llama3-8B-IT and Gemma2-9B-IT.
-Metrics include model LC win rate on AlpacaEval2 and raw win rate on Arena-Hard.
-Fitted dotted curves apply scaling laws from Gao et al.
-(2023) to direct alignment, with GPT-4 win rates substituting gold reward model scores.
+Figure 6: Over-optimization patterns for RePO on Llama3-8B-IT and Gemma2-9B-IT, using AlpacaEval2 LC win rates and Arena-Hard raw win rates.
+Dotted curves represent theoretical fits based on Gao et al.
+[9]’s scaling laws, using GPT-4 win rates instead of standard reward model scores.
 Scaling Law Fits.
-Previous work (Gao et al., 2023; Rafailov et al., 2024) has established scaling laws for reward model scores as a function of the KL divergence between the initial and optimized policies.
+Previous work [9, 16] has established scaling laws for reward model scores as a function of the KL divergence between the initial and optimized policies.
 In contrast, we eliminate the reference model and the associated computational cost of calculating KL divergence.
@@
 Interestingly, we find that this scaling law provides an accurate relationship between $d$ and win rates for RePO.
-3.5 RePO++
-Building on the previous discussion, we identify the advantages of RePO in finding a “cutoff point” to control over-optimization.
-However, a limitation persists: for cases where the implicit reward margin is smaller than $\gamma$, their gradient weights become unit, failing to differentiate between samples.
-One potential solution is to apply a non-linear activation function to weight the gradient of each pair. Specifically, we extend RePO to DPO and SimPO, introducing a new method, RePO++.
+3.4 RePO++: Exploring Extensions of ReLU-based Filtering
+While exploring ReLU’s thresholding behavior, we observed an interesting limitation: for cases where the implicit reward margin is smaller than $\gamma$, their gradient weights become uniform, not differentiating between samples of varying difficulty.
+This observation naturally led us to wonder: could we preserve the effective filtering mechanism while reintroducing some degree of weighting?
+To explore this question, we experimented with combining ReLU’s binary filtering with SimPO’s continuous weighting:
 $\mathcal{L}_{\text{RePO++}}(\pi_{\theta})=-\mathbb{E}_{\mathcal{D}}\left[\log\sigma\left(-\text{ReLU}\left(-\beta\left(M_{\theta}-\gamma\right)\right)\right)\right],$ (12)
-What does the RePO++ update do? We again consider analyzing the gradient of the loss function RePO++ for an intuitive understanding. The gradient w.r.t. the parameters $\theta$ is given by:
+This exploration was a natural follow-up to our main discovery about ReLU’s effectiveness, rather than our primary contribution.
+We were curious to see whether combining the best aspects of both approaches might yield additional insights about preference learning mechanisms.
+What does this combined approach reveal? To understand the behavior of this extension, we examined its gradient with respect to the parameters $\theta$:
 $\nabla_{\theta}\mathcal{L}_{\mathrm{RePO++}}(\pi_{\theta})=-\beta\mathbb{E}_{\mathcal{D}}\left[s_{\theta}\cdot\mathbb{I}_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right],$ (13)
 where $s_{\theta}=\sigma\left(\beta\left(-M_{\theta}+\gamma\right)\right)$ and $\mathbb{I}_{\theta}$ is an indicator function that is 1 if $M_{\theta}<\gamma$ and 0 otherwise.
-Intuitively, This gradient combines the strengths of SimPO and RePO: it scales updates by $s_{\theta}$ and filters them using $\mathbb{I}_{\theta}$, ensuring that the model focuses on less-separated pairs, with smaller separations receiving higher weights.
+We observed that this gradient combines properties we discovered in both approaches: it scales updates by $s_{\theta}$ (similar to SimPO) and filters them using $\mathbb{I}_{\theta}$ (the key discovery in our ReLU exploration), focusing the model on less-separated pairs while giving higher weights to smaller separations.
 Adaptation with RePO++.
@@
 For instance, as shown in Equation 12, replacing $M_{\theta}$ with $\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}$ seamlessly integrates RePO++ into DPO, forming a ReLU-enhanced version of DPO.
-4 Theoretical Analysis of RePO
-Next, we relate preference optimization to a supervised binary classification task and show how RePO leverages the convex envelope of the 0-1 loss to achieve both theoretical optimality and efficient optimization.
+4 Theoretical Analysis: ReLU’s Optimality in Preference Learning
+Next, we establish a surprising theoretical connection between preference optimization and binary classification, revealing why our simple ReLU-based approach achieves superior performance.
 Following Tang et al.
-(2024), we reformulate reward modeling as a supervised binary classification task.
-Given data pairs $(z,l)$ with $z\in\mathbb{R}^{k}$ and labels $l\in\{-1,1\}$, the goal is to learn a predictor ^$\hat{\ell}(z)\in\mathbb{R}$ such that $\mathrm{sign}(\hat{\ell}(z))$ matches $l$.
-The classification accuracy is given by: $\frac{1}{2}\mathbb{E}[\mathrm{sign}(\hat{\ell}(z)\cdot l)]+\frac{1}{2},$ and the corresponding loss function is:
-$\mathcal{L}_{0\text{-}1}(\hat{\ell}):=\mathbb{E}\left[1-\mathrm{sign}(\hat{\ell}(z)\cdot l)\right].$
-The above loss, known as the 0-1 loss is non-convex. We therefore approximate it using convex functions $f:\mathbb{R}\to\mathbb{R}$:
-$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(\hat{\ell}(z)\cdot l)\right].$
-In the context of reward modeling, given response pairs $(y_{w},y_{l})$ with preference $y_{w}\succ y_{l}$, we construct binary classification samples with $\ell=1$.
-For a pointwise reward model $r_{\phi}:\mathcal{Y}\to\mathbb{R}$, the natural parameterization is: ^$\hat{\ell}(y_{w},y_{l})=r_{\phi}(y_{w})-r_{\phi}(y_{l}),$ leading to the loss:
-$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(r_{\phi}(y_{w})-r_{\phi}(y_{l}))\right].$
-Different choices of $f(x)$ correspond to distinct loss functions.
-Specifically, when $f(x)=\mathbb{I}(x<0)$, the 0-1 loss is obtained; when $f(x)=-\log\sigma(x)$, we obtain the logistic loss used in SimPO; and when $f(x)=\text{ReLU}(-x)$, it corresponds to the loss function used in RePO.
+[17], preference learning can be reformulated as binary classification.
+Given pairs $(z,l)$ where $z\in\mathbb{R}^{k}$ and $l\in\{-1,1\}$, we aim to learn a predictor $\hat{\ell}(z)$ whose sign matches $l$.
+The classification accuracy is: $\frac{1}{2}\mathbb{E}[\mathrm{sign}(\hat{\ell}(z)\cdot l)]+\frac{1}{2}.$ This corresponds to minimizing the 0-1 loss:
+$\mathcal{L}_{0\text{-}1}(\hat{\ell}):=\mathbb{E}\left[1-\mathrm{sign}(\hat{\ell}(z)\cdot l)\right]$ (14)
+For preference data $(y_{w},y_{l})$ where $y_{w}\succ y_{l}$, we set $l=1$ and parameterize $\hat{\ell}(y_{w},y_{l})=r_{\phi}(y_{w})-r_{\phi}(y_{l})$, yielding the objective:
+$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(r_{\phi}(y_{w})-r_{\phi}(y_{l}))\right]$ (15)
+Where $f$ determines the surrogate loss: $f(x)=\mathbb{I}(x<0)$ gives the 0-1 loss, $f(x)=-\log\sigma(x)$ yields SimPO’s logistic loss, and $f(x)=\text{ReLU}(-x)$ gives our method’s loss.
+Our key insight comes from analyzing the convex envelope of the 0-1 loss:
 Definition 4.1.
-Let $D\subseteq\mathbb{R}$ be a closed convex set. The convex envelope of $\mathcal{L}_{0\text{-}1}$ over $D$ is:
-$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)$
-::=:= $\sup\left\{h(x)\mid h\text{ is convex},\,h\leq\mathcal{L}_{0\text{-}1}\,\forall x\in D\right\}.$ (17)
-Theorem 4.2 (Convex Envelope of 0-1 Loss).
-For $D=[-a,b]$ with $a,b>0$, the convex envelope of $\mathcal{L}_{0\text{-}1}(x)=\mathbb{I}(x<0)$ is:
-$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)=\frac{1}{a}\mathrm{ReLU}(-x).$
+The convex envelope of $\mathcal{L}_{0\text{-}1}$ over a closed convex set $D\subseteq\mathbb{R}$ is:
+$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x):=\sup\left\{h(x)\mid h\text{ is convex},\,h\leq\mathcal{L}_{0\text{-}1}\,\forall x\in D\right\}$ (16)
+Theorem 4.2 (ReLU as Convex Envelope).
+For D=[a,b]D=[-a,b] with $a,b>0$, the convex envelope of $\mathcal{L}_{0\text{-}1}(x)=\mathbb{I}(x<0)$ is:
+$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)=\frac{1}{a}\mathrm{ReLU}(-x)$ (17)
+This remarkable result reveals that ReLU provides the tightest possible convex approximation to the ideal 0-1 loss, explaining its empirical effectiveness. Furthermore:
 Corollary 4.3 (Optimality Preservation).
 Let $D\subseteq\mathbb{R}$ be convex. Then:
-  1. 1.
-    For binary classification with loss $\mathcal{L}_{0\text{-}1}$:
-    $\arg\min_{\hat{\ell}}\mathcal{L}_{0\text{-}1}(\hat{\ell})=\arg\min_{\hat{\ell}}\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(\hat{\ell}).$
-  2. 2.
-    When $D=[-a,b]$:
-    $\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x).$
-Remark 4.4.
-Corollary 4.3 establishes: (i) equivalence between minimizing the 0-1 loss and its convex envelope, (ii) exact optimality preservation for the ReLU surrogate.
-This guarantees gradient-based optimization converges to solutions matching the theoretical optimum of the intractable 0-1 loss.
-Corollary 4.5 (Logistic Loss is Not the Convex Envelope of 0-1 Loss).
+[…]
+And for D=[a,b]D=[-a,b]:
+$\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x)$ (19)
+This guarantees that gradient-based optimization of our ReLU surrogate converges to solutions matching the theoretical optimum of the intractable 0-1 loss. Importantly:
+Corollary 4.4 (Logistic Loss Suboptimality).
 The logistic loss $f_{\mathrm{log-sigmoid}}(x)=-\log\sigma(x)$ is not the convex envelope of $\mathcal{L}_{0\text{-}1}$.
-Remark 4.6.
-Unlike SimPO’s logistic loss (Corollary 4.5), the ReLU-based surrogate inherits the 0-1 loss’s global optimality.
-This explains RePO’s robustness to suboptimal local minima during gradient-based training.
+This theoretical foundation explains why our simple ReLU-based approach consistently outperforms more complex mechanisms like SimPO’s sigmoid weighting — ReLU provides optimality guarantees that logistic loss cannot match, while being computationally more efficient.
 5 Experiments
-In this section, we present our main experimental findings, demonstrating that RePO consistently achieves better results than existing preference optimization methods across multiple benchmarks.
+In this section, we examine how our simplified ReLU-based approach behaves across different models and settings.
+Rather than focusing solely on performance gains, we explore patterns that help explain why such a simple mechanism works effectively in practice.
 5.1 Experimental Setup
-The main experimental setup has been described in Section 3.2.
-We increase a model: Mistral2-7B (Jiang et al., 2023).
-Additionally, the v0.2 Llama3-Instruct setup uses RLHFlow/ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024b) as the reward model for ranking generated data, significantly enhancing performance.
-As for baselines, we compare RePO with several state-of-the-art preference optimization methods: DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), IPO (Azar et al., 2023), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and R-DPO (Park et al., 2024).
-We also include the SFT model as a baseline.
-We thoroughly tune the hyperparameters for each baseline and report the best performance.
-Further details can be found in Appendix C.1.
-We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks (Beeching et al., 2023), with additional details in in Appendix C.2.
+The core experimental configuration extends our investigation from Section 3.2 to include Mistral2-7B [18] alongside previously examined models.
+For the Llama3-Instruct v0.2 experiments, we employed the RLHFlow/ArmoRM-Llama3-8B-v0.1 [19] reward model for ranking generated data.
+We benchmark our approach against established preference optimization methods: DPO [6], SimPO [11], IPO [20], CPO [21], KTO [22], ORPO [23], and R-DPO [24], with SFT models serving as baselines.
+Implementation details are provided in Appendix D.1.
+We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks [25], with additional details in in Appendix D.2.
 The code is available at https://github.com/junkangwu/RePO.
@@
 AE 2 AH AE 2 AH AE 2 AH AE 2 AH
-LC (%) WR (%) WR (%) LC (%) WR (%) WR (%) LC (%) WR (%) WR (%) LC (%) WR (%) WR (%)
+LC WR WR LC WR WR LC WR WR LC WR WR
 SFT 24.0 23.6 22.4 19.0 15.4 12.9 24.0 23.6 22.4 48.7 36.5 42.1
@@
 RePO 46.7 41.1 33.3 30.4 33.6 20.3 57.7 51.1 35.2 73.6 66.6 59.1
+5.2 Result Comparisons
+Observation: Simple ReLU thresholding exhibits surprising effectiveness.
+Table 1 reveals an unexpected pattern: despite removing components previously thought essential, the simple ReLU-based approach consistently performs well across all evaluated models and benchmarks.
+This finding aligns with our theoretical analysis showing that binary thresholding directly approximates the convex envelope of the 0-1 loss.
+On AlpacaEval 2, we observe improvements of 0.2-2.8 points in LC win rates across different configurations compared to the strongest baselines.
+5.3 Methodology Comparisons
+Beyond alignment, we also compare the methodologies of these preference learning methods. Our method plays a hub to connect these methods.
+Relation to SimPO.
+SimPO employs sigmoid weighting via log-sigmoid activation to attenuate gradients as reward margins increase, mitigating over-optimization.
+RePO can be viewed as “SimPO without log-sigmoid,” replacing this continuous scaling with binary filtering.
+To validate this relationship, we integrated a ReLU-based filtering mechanism into SimPO (cf.
+RePO++ Equation 12).
+Table 2 confirms that ReLU’s filtering mechanism enhances performance.
+As demonstrated in Section 3.4, RePO++ directly addresses over-optimization while retaining the benefits of some weighting.
+Relation to SLiC-HF.
+RePO can be characterized as “SLiC-HF without SFT regularization”.
+To ensure a fair comparison (while disregarding differences in length normalization), we investigated the impact of SFT regularization by varying its coefficient, $\lambda$.
+The results, presented in Appendix Table 6 and further details in Appendix D.5, indicate that this additional regularization term offers no discernible improvement.
+This suggests that SFT regularization targets a different optimization challenge, distinct from the direct over-optimization problem RePO addresses.
+Relation to DPO.
+Mathematically, DPO is equivalent to SimPO when the margin $\gamma$ is defined as $\log\pi_{\text{ref}}(y_{w}\mid x)-\log\pi_{\text{ref}}(y_{l}\mid x)$ (ignoring length normalization).
+However, directly substituting the log-sigmoid function with ReLU in DPO’s formulation leads to a significant performance degradation (see Appendix Table 7 and Appendix D.6).
+This underscores the critical role of the threshold $\gamma$ in determining the effectiveness of over-optimization prevention.
+As identified by Wu et al.
+[26], reference model based reward margins are often unreliable as target margins, which explains why SimPO’s explicit $\gamma$ parameter is effective for preference learning.
+5.4 Effect of ReLU Filtering Across Methods
+Having observed the effectiveness of binary thresholding, we naturally questioned whether this mechanism might enhance other preference learning approaches.
+Table 2 shows that integrating ReLU filtering consistently improved performance across both DPO and SimPO frameworks, suggesting that selective gradient application based on margin thresholds provides benefits beyond our specific implementation.
+Our experiments with the combined approach (cf.
+RePO++ in Section 3.4) revealed particularly strong improvements when applied to DPO (5%–12% gains), with notable performance on Arena-Hard (reaching 65.7).
+This design effectively mitigates over-optimization while preserving the benefits of the original scheme.
 Table 2: Performance improvements of RePO and RePO++ over DPO and SimPO.
@@
 AE 2 AH AE 2 AH
-LC (%) WR (%) WR (%) LC (%) WR (%) WR (%)
+LC WR WR LC WR WR
 DPO 48.2 47.5 35.2 70.4 66.9 58.8
 w.
-RePO $50.3^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.4\%}$ $51.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.1\%}$ $38.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+8.5\%}$ $73.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.8\%}$ $71.0^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.1\%}$ $64.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.2\%}$
+RePO $50.3^{\color[rgb]{0,0,0}+4.4\%}$ $51.8^{\color[rgb]{0,0,0}+9.1\%}$ $38.2^{\color[rgb]{0,0,0}+8.5\%}$ $73.8^{\color[rgb]{0,0,0}+4.8\%}$ $71.0^{\color[rgb]{0,0,0}+6.1\%}$ $64.2^{\color[rgb]{0,0,0}+9.2\%}$
 w.
-RePO++ $50.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+5.4\%}$ $52.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.9\%}$ $37.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+5.7\%}$ $71.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.0\%}$ $69.5^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.9\%}$ $65.7^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+11.7\%}$
+RePO++ $50.8^{\color[rgb]{0,0,0}+5.4\%}$ $52.2^{\color[rgb]{0,0,0}+9.9\%}$ $37.2^{\color[rgb]{0,0,0}+5.7\%}$ $71.8^{\color[rgb]{0,0,0}+2.0\%}$ $69.5^{\color[rgb]{0,0,0}+3.9\%}$ $65.7^{\color[rgb]{0,0,0}+11.7\%}$
 SimPO $55.6$ $49.6$ $33.6$ $72.4$ $65.0$ $57.8$
 w.
-RePO $57.7^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.8\%}$ $51.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.0\%}$ $35.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.8\%}$ $73.6^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+1.7\%}$ $66.6^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.5\%}$ $59.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.2\%}$
+RePO $57.7^{\color[rgb]{0,0,0}+3.8\%}$ $51.1^{\color[rgb]{0,0,0}+3.0\%}$ $35.2^{\color[rgb]{0,0,0}+4.8\%}$ $73.6^{\color[rgb]{0,0,0}+1.7\%}$ $66.6^{\color[rgb]{0,0,0}+2.5\%}$ $59.1^{\color[rgb]{0,0,0}+2.2\%}$
 w.
-RePO++ $56.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+0.9\%}$ $50.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+1.0\%}$ $35.9^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.8\%}$ $74.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.3\%}$ $66.5^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.3\%}$ $59.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.5\%}$
-5.2 Main Results
-RePO consistently and significantly outperforms existing preference optimization methods.
-As shown in Table 1, all preference optimization methods outperform the SFT model.
-Nevertheless, RePO — despite its simplicity — achieves the strongest overall performance across all benchmarks and settings.
-The consistent gains highlight its robustness and efficiency.
-Specifically, RePO surpasses the best-performing baseline by 0.2 to 2.8 points on the AlpacaEval 2 LC win rate under various configurations.
-On Arena-Hard, RePO generally performs better than competing methods, although CPO achieves slightly higher scores at times.
-We observe that CPO tends to produce responses that are approximately 50% longer.
-Because Arena-Hard does not penalize response length, longer outputs may occasionally yield higher evaluations.
-5.3 Adaptations of RePO
-RePO is a versatile approach that can be adapted to different optimization settings.
-In this work, we explore two adaptations: RePO and RePO++, applied to both DPO and SimPO.
-The detailed formulations are provided in Section 3.5.
-As shown in Table 2, integrating RePO with DPO and SimPO consistently enhances performance.
-Notably, RePO++ achieves superior results in most scenarios, which we attribute to its combination of ReLU activation and the original weighting function $s_{\theta}$.
-This design effectively mitigates overoptimization while preserving the benefits of the original scheme.
-Interestingly, applying RePO and RePO++ to DPO yields particularly significant improvements (5%–12%), achieving a high score of 65.7 on the Arena-Hard benchmark.
-5.4 Dynamic Margin Scheduling
-To further investigate the impact of the target reward margin $\gamma$ , we conduct an ablation study using dynamic values for $\gamma$ Specifically, we implement two strategies: (i) $\gamma$ increases from small to large, and (ii) $\gamma$ decreases from large to small.
-Using RePO, we control the initial and final values of $\gamma$ , adjusting it uniformly over the course of one training epoch.
-As shown in Figure 7, both excessively small and large values of $\gamma$ lead to suboptimal performance (e.g., , $1.0\rightarrow 0.8$ , $0.0\rightarrow 0.2$ ).
-In contrast, starting with a moderately large value of $\gamma$ and gradually decreasing it during training (e.g., , $1.0\rightarrow 0.2$ ) results in improved model performance.
-We interpret this behavior as follows: in the early stages of training, the model is underfitting and thus benefits from stronger updates.
-In later stages, reducing $\gamma$ helps mitigate overfitting, effectively addressing this issue.
-Figure 7: Performance of RePO with dynamic $\gamma$. The dashed line represents RePO with a fixed $\gamma$.
+RePO++ $56.1^{\color[rgb]{0,0,0}+0.9\%}$ $50.1^{\color[rgb]{0,0,0}+1.0\%}$ $35.9^{\color[rgb]{0,0,0}+6.8\%}$ $74.1^{\color[rgb]{0,0,0}+2.3\%}$ $66.5^{\color[rgb]{0,0,0}+2.3\%}$ $59.8^{\color[rgb]{0,0,0}+3.5\%}$
+5.5 Dynamic Margin Scheduling and Curriculum Learning
+Our investigation into the role of target reward margin $\gamma$ led us to an unexpected discovery about curriculum learning.
+We experimented with dynamic scheduling of $\gamma$ throughout training, implementing two strategies: (i) increasing $\gamma$ from small to large, and (ii) decreasing $\gamma$ from large to small.
+Figure 7 reveals a striking pattern: starting with a moderately large value of $\gamma$ and gradually decreasing it ($1.0\rightarrow 0.2$) naturally creates an effective curriculum that improves model performance.
+In contrast, both excessively large values ($1.0\rightarrow 0.8$) and small values ($0.0\rightarrow 0.2$) led to suboptimal outcomes.
+Figure 7: Exploration of dynamic $\gamma$ scheduling on Llama3-Instruct v0.2 (8B).
+The dashed line represents performance with a fixed $\gamma$.
+We observed that decreasing $\gamma$ from an initially larger value creates a natural curriculum that enhances performance.
+This observation reveals an intriguing self-regulating property: early in training when the model is underfitting, a larger $\gamma$ permits more aggressive updates across more examples.
+As training progresses, the decreasing $\gamma$ naturally focuses learning on increasingly challenging examples, effectively preventing over-optimization.
+This emergent curriculum behavior, arising from a simple parameter schedule, suggests that binary thresholding captures fundamental learning dynamics that more complex mechanisms might obscure.
 6 Discussion
-Conclusion.
-In this work, we propose RePO, a simple preference optimization by replacing logistic weighting with a ReLU-based max-margin loss, eliminating the need for hyperparameter $\beta$.
-Theoretically, RePO is characterized as SimPO’s $\beta\rightarrow\infty$ limit, collapsing to binary thresholding for data filtering.
-Empirically, it matches or outperforms DPO and SimPO on AlpacaEval 2 and Arena-Hard with only a single tunable parameter ($\gamma$).
-Its gradient dynamics prioritize challenging samples, mitigating over-optimization while maintaining training stability.
-This work establishes RePO as a robust, hyperparameter-efficient alternative for offline alignment.
-Limitations and Future Work.
-One limitation of our approach is that it currently operates in an offline setting, and future research will focus on extending it to online reinforcement learning frameworks.
-Integrating repo with online RL techniques could enhance real-time adaptability and scalability.
-Another direction for future work is to explore how the cutoff point strategy can be further refined to sustain performance improvements in self-play scenarios (Chen et al., 2024b), particularly in highly dynamic environments.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Conclusion Our exploration of simple ReLU activation in preference learning has revealed several key insights.
+We found that binary thresholding, implemented through a straightforward ReLU function, provides an effective mechanism for preventing over-optimization in language model alignment.
+Our theoretical analysis showed that this seemingly simple approach is, in fact, the convex envelope of the ideal 0-1 loss function, explaining its surprising effectiveness.
+Rather than developing yet another complex preference optimization method, our work uncovered how fundamental properties like data selection and implicit curriculum learning emerge naturally from basic principles.
+Limitations and future directions.
+Our current exploration is limited to offline preference learning settings.
+Future work could investigate how these insights might extend to online learning scenarios, where preferences are gathered interactively.
+Additionally, while we found that a fixed margin threshold works well in practice, exploring adaptive or context-aware thresholds might further improve performance in highly dynamic environments.
+The relationship between binary filtering and self-play scenarios [27] — where the model generates its own feedback — is another promising direction that could lead to more scalable alignment techniques.
+Beyond alignment, our work connects to LLM reasoning research [28, 29].
+Future work should investigate how KL penalties and gradient clipping in GRPO [30] and PPO [5] balance preventing over-optimization against preserving reasoning capabilities — a critical consideration for advancing alignment methodologies.
+Acknowledgments and Disclosure of Funding
+This research was supported by the National Science and Technology Major Project (2023ZD0121102), the National Natural Science Foundation of China (U24B20180, 62302321), and the Fundamental Research Funds for the Central Universities (WK2100250065).
+This research also benefited from the advanced computing resources provided by the Supercomputing Center of the USTC.
 Appendix A Related Works
 Reinforcement learning from human feedback.
-RLHF is a technique designed to align large language models with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Azar et al., 2023).
-Traditional RLHF is typically structured in three stages: supervised fine-tuning (Zhou et al., 2023; Taori et al., 2023; Geng et al., 2023; Conover et al., 2023; Köpf et al., 2023; Ding et al., 2023), reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024a; Lightman et al., 2023; Havrilla et al., 2024; Lambert et al., 2024), and policy optimization (Schulman et al., 2017; Anthony et al., 2017).
+RLHF is a technique designed to align large language models with human preferences and values [31, 32, 33, 20].
+Traditional RLHF is typically structured in three stages: supervised fine-tuning [34, 35, 36, 37, 38, 39], reward modeling [9, 40, 41, 42, 43, 44], and policy optimization [5, 45].
 In the third stage, Proximal Policy Optimization (PPO) is widely adopted.
-In contrast, RLOO (Ahmadian et al., 2024) reduces the GPU memory footprint of RLHF by eliminating the Critic model and leverages a Leave-One-Out strategy to achieve superior performance.
-GRPO (Shao et al., 2024), another variant of PPO, improves mathematical reasoning abilities while optimizing memory usage by replacing the Leave-One-Out method with a direct subtraction of the mean of all samples for a given prompt.
+In contrast, RLOO [46] reduces the GPU memory footprint of RLHF by eliminating the Critic model and leverages a Leave-One-Out strategy to achieve superior performance.
+GRPO [47], another variant of PPO, improves mathematical reasoning abilities while optimizing memory usage by replacing the Leave-One-Out method with a direct subtraction of the mean of all samples for a given prompt.
 Offline preference optimization.
 In addition to DPO, several alternative preference optimization objectives have been proposed.
-IPO (Azar et al., 2023) addresses overfitting issues inherent in DPO.
-ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) aim to remove reliance on a reference model.
-R-DPO (Park et al., 2024) targets the reduction of exploitation due to sequence length, while KTO (Ethayarajh et al., 2024) handles preference optimization in the absence of pairwise data.
-CPO (Xu et al., 2024) and $\beta$-DPO (Wu et al., 2024b) focus on improving the quality of preference data.
+IPO [20] addresses overfitting issues inherent in DPO.
+ORPO [23] and SimPO [11] aim to remove reliance on a reference model.
+R-DPO [24] targets the reduction of exploitation due to sequence length, while KTO [22] handles preference optimization in the absence of pairwise data.
+CPO [21] and $\beta$-DPO [48] focus on improving the quality of preference data.
 Another research direction addresses noise in offline alignment, which arises from the need to construct pairwise data.
-rDPO (Chowdhury et al., 2024), a variant of DPO, mitigates preference noise and enhances policy robustness, while DrDPO (Wu et al., 2024a) applies distributed robust optimization to tackle this issue.
-Other works have approached the problem through divergence regularization (Wang et al., 2024a; Huang et al., 2024), selection of high-quality data (Khaki et al., 2024; Ye et al., 2024), or reweighting loss functions (Wang et al., 2024c; Muldrew et al., 2024; Yang et al., 2024).
+rDPO [49], a variant of DPO, mitigates preference noise and enhances policy robustness, while DrDPO [50] applies distributed robust optimization to tackle this issue.
+Other works have approached the problem through divergence regularization [51, 52], selection of high-quality data [53, 54], or reweighting loss functions [55, 56, 57].
 Iterative Preference Optimization.
@@
 To address this, iterative preference optimization techniques have been proposed.
-These methods iteratively update the reference model using the most recent policy model or generate new preference pairs in each iteration (Dong et al., 2024; Kim et al., 2024; Rosset et al., 2024; Xiong et al., 2024; Yuan et al., 2024; Chen et al., 2024b; Wu et al., 2024c; Gao et al., 2024).
-For instance, SPIN (Chen et al., 2024b) employs a self-play framework to fine-tune the model in a supervised manner, while Yuan et al.
-(2024) annotate preferences throughout the iterative process.
+These methods iteratively update the reference model using the most recent policy model or generate new preference pairs in each iteration [58, 59, 60, 61, 62, 27, 63, 64].
+For instance, SPIN [27] employs a self-play framework to fine-tune the model in a supervised manner, while Yuan et al.
+[62] annotate preferences throughout the iterative process.
 REBEL improves sample quality by regressing the relative reward.
-Additionally, (Calandriello et al., 2024) generates data using a mixture policy, similar to the Nash-MD algorithm (Rosset et al., 2024).
-Appendix B Proofs
-B.1 Proof of Lemma 3.1
+Additionally, [65] generates data using a mixture policy, similar to the Nash-MD algorithm [60].
+Appendix B Broader Impacts
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here
+Appendix C Proofs
+C.1 Proof of Lemma 3.1
 See 3.1
@@
 Case 1: $M_{\theta}<\gamma$
-Here, $-M_{\theta}+\gamma>0$. As $\beta\to\infty$,
+Here, $M_{\theta}+\gamma>0$. As $\beta\to\infty$,
 $\lim_{\beta\to\infty}\sigma\big(\beta(-M_{\theta}+\gamma)\big)=\lim_{z\to\infty}\sigma(z)=1=\mathbb{I}(M_{\theta}<\gamma).$
 Case 2: $M_{\theta}>\gamma$
-Here, $-M_{\theta}+\gamma<0$. As $\beta\to\infty$,
+Here, $M_{\theta}+\gamma<0$. As $\beta\to\infty$,
 $\lim_{\beta\to\infty}\sigma\big(\beta(-M_{\theta}+\gamma)\big)=\lim_{z\to-\infty}\sigma(z)=0=\mathbb{I}(M_{\theta}<\gamma).$
@@
 $\lim_{\beta\to\infty}\nabla_{\theta}\mathcal{L}_{\mathrm{SimPO}}$ $=-\lim_{\beta\to\infty}\beta\mathbb{E}_{\mathcal{D}}\left[s_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]$ (22)
-[…]
+$=-\mathbb{E}_{\mathcal{D}}\left[\lim_{\beta\to\infty}\beta s_{\theta}\cdot\left(\nabla_{\theta,y_{w}}-\nabla_{\theta,y_{l}}\right)\right]\quad\text{(Dominated Convergence Theorem)}$ (23)
 To resolve the $\beta$ scaling, observe that for $M_{\theta}\neq\gamma$:
@@
 which completes the proof. ∎
-B.2 Proof of Theorem 4.2
+C.2 Proof of Theorem 4.2
 See 4.2
@@
 Epigraph Characterization: The epigraph $\mathrm{epi}(h)$ coincides with the convex hull of $\mathrm{epi}(\mathcal{L}_{0\text{-}1})\cap(D\times\mathbb{R})$.
-The affine segment $h(x)=-\frac{x}{a}$ on $[-a,0)$ connects the points $(-a,1)$ and $(0,0)$, forming the tightest convex fit to the 0-1 loss’s discontinuity.
-By Theorem 1 in (Princeton University, Accessed 2025), this construction achieves the convex envelope.
+The affine segment $h(x)=-\frac{x}{a}$ on [a,0)[-a,0) connects the points (a,1)(-a,1) and $(0,0)$, forming the tightest convex fit to the 0-1 loss’s discontinuity.
+By Theorem 1 in [66], this construction achieves the convex envelope.
 ∎
-B.3 Proof of Corollary 4.3
+C.3 Proof of Corollary 4.3
 See 4.3
@@
 Since $\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x^{*})\leq\mathcal{L}_{0\text{-}1}(x^{*})$ and $\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}$ attains its minimum at the same points as $\mathcal{L}_{0\text{-}1}$, $x^{*}$ must also minimize $\mathcal{L}_{0\text{-}1}$.
-Part 2: For $D=[-a,b]$, both $\mathcal{L}_{0\text{-}1}(x)$ and $\frac{1}{a}\mathrm{ReLU}(-x)$ attain their minimum value 0 on $[0,b]$.
+Part 2: For D=[a,b]D=[-a,b], both $\mathcal{L}_{0\text{-}1}(x)$ and $\frac{1}{a}\mathrm{ReLU}(-x)$ attain their minimum value 0 on $[0,b]$.
 For $x\in[-a,0)$, $\frac{1}{a}\mathrm{ReLU}(-x)$ is strictly decreasing, achieving its minimum at $x=0$.
@@
 $\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x)=[0,b].$
-B.4 Proof of Corollary 4.5
-See 4.5
+C.4 Proof of Corollary 4.4
+See 4.4
 Proof.
-We demonstrate violation of the convex envelope’s defining property. Consider $D=[-1,1]$:
+We demonstrate violation of the convex envelope’s defining property. Consider D=[1,1]D=[-1,1]:
 1. Underestimation Failure: For $x>0$:
-$-\log\sigma(x)=-\log\left(\frac{1}{1+e^{-x}}\right)=\log(1+e^{-x})>0=\mathcal{L}_{0\text{-}1}(x)$
+$\log\sigma(x)=-\log\left(\frac{1}{1+e^{-x}}\right)=\log(1+e^{-x})>0=\mathcal{L}_{0\text{-}1}(x)$
 Thus $f_{\mathrm{log-sigmoid}}\not\leq\mathcal{L}_{0\text{-}1}$ over $D$, violating the envelope requirement.
@@
 Hence $f_{\mathrm{log-sigmoid}}$ cannot be the convex envelope. ∎
-Appendix C Experiments
-C.1 Implementation Details
+Appendix D Experiments
+D.1 Implementation Details
 Empirical observations indicate significant performance sensitivity to model parameter initialization and learning rate selection across compared methods.
@@
   • Warmup: 10% of total training steps
-  • Optimizer: Adam (Kingma & Ba, 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
+  • Optimizer: Adam [13] ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
   • Sequence length: 2048 tokens (fixed for all inputs)
@@
 Method Objective Hyperparameter
-DPO (Rafailov et al., 2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in[0.01,0.05,0.1]$
-IPO (Azar et al., 2023) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
-CPO (Xu et al., 2024) $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
-KTO (Ethayarajh et al., 2024) $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$
+SLiC-HF [7] $\max\left(0,\delta-\log\pi_{\theta}(y_{w}|x)+\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda\in[0.1,0.5,1.0,10.0]$
+$\delta\in[0.1,0.5,1.0,2.0]$
+DPO [6] $\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in[0.01,0.05,0.1]$
+IPO [20] $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
+CPO [21] $\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
+KTO [22] $\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$
 $\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\text{KL}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)\right]$ $\beta\in[0.01,0.05,0.1]$
-ORPO (Hong et al., 2024) $-\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$ $\lambda\in[0.1,0.5,1.0,2.0]$
+ORPO [23] $\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$ $\lambda\in[0.1,0.5,1.0,2.0]$
 $\text{where}\,\,p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$
-R-DPO (Park et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
+R-DPO [24] $\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
 $\beta\in[0.01,0.05,0.1]$
-SimPO (Meng et al., 2024) $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in[2.0,4.0,6.0,8.0]$
+SimPO [11] $\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in[2.0,4.0,6.0,8.0]$
 $\gamma\in[0.3,0.5,1.0,1.2,1.4,1.6]$
-RePO $\text{ReLU}[-(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma)]$ $\gamma\in[0.2,0.4,0.6,0.8]$
+RePO $\text{ReLU}[-(\frac{1}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma)]$ $\gamma\in[0.2,0.4,0.5,0.6,0.8]$
 Table 4: The hyperparameter values in RePO used for each training setting.
@@
 The experimental setup follows the guidelines provided in the alignment-handbook repository [^3] https://github.com/huggingface/alignment-handbook, ensuring reproducibility and consistency with established practices.
-C.2 Downstream Task Evaluation
+D.2 Downstream Task Evaluation
 Table 5: Downstream task evaluation results of tasks on the huggingface open leaderboard.
@@
 RePO 65.00 68.09 80.50 64.38 76.16 69.37 70.58
-To assess the impact of RePO on downstream task performance, we evaluate models trained with different preference optimization methods on a diverse set of tasks from the Huggingface Open Leaderboard (Beeching et al., 2023).
-The tasks include MMLU (Hendrycks et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2022), Winograd (Levesque et al., 2012), and GSM8K (Cobbe et al., 2021).
-We adhere to standard evaluation protocols and present the results for all models in Table 5.
+To assess the impact of RePO on downstream task performance, we evaluate models trained with different preference optimization methods on a diverse set of tasks from the Huggingface Open Leaderboard [25].
+The tasks include MMLU [67], ARC [68], HellaSwag [69], TruthfulQA [70], Winograd [71], and GSM8K [72].
+We adhere to standard evaluation protocols and present the results for all models in Table˜5.
 Overall Performance.
@@
 This suggests that RePO has strengths in some domains, but it may benefit from further refinement to improve performance across all tasks.
-Conclusion and Future Work.
-In summary, RePO shows promising results across a range of downstream tasks, with notable improvements in tasks related to commonsense reasoning and truthfulness.
-However, it faces challenges in preserving general knowledge and mathematical reasoning.
-Future work could focus on incorporating additional strategies to mitigate performance drops in these areas, possibly through the integration of task-specific objectives or enhanced fine-tuning techniques.
-Additionally, further empirical studies are needed to explore the impact of different preference optimization datasets and objectives on downstream performance.
-C.3 RePO with varying $\gamma$
+D.3 RePO with varying $\gamma$
 Figure 8: Impact of the hyperparameter $\gamma$ on LC Win Rate and Raw Win Rate for Llama3-8B-Instruct (left) and Gemma2-9B-Instruct (right).
@@
 Future work could explore adaptive strategies for dynamically tuning $\gamma$, ensuring that preference optimization enhances alignment without compromising generalization.
-C.4 Analsis on RePO++
+D.4 Analsis on RePO++
 Figure 9 illustrates the relationship between the gradient and the implicit reward margin.
@@
 Figure 9: Gradient weighting functions of RePO++ ($s_{\theta}\cdot\mathbb{I}(M_{\theta}<\gamma)$).
+D.5 Analysis of the Relationship Between SLiC-HF and RePO
+To provide deeper insights into the relationship between SLiC-HF and RePO, we conducted additional experiments examining the effect of SFT regularization—a core component of SLiC-HF that is absent in our method.
+As demonstrated in Table 6, we systematically evaluated performance across varying values of the regularization coefficient $\lambda$.
+Mathematical Comparison.
+From a formulation perspective, SLiC-HF combines a hinge loss term with an SFT regularization component that penalizes deviation from the reference model.
+Specifically, the SFT regularization is controlled by parameter $\lambda$, which balances preference optimization against model drift.
+By contrast, RePO eliminates this regularization entirely, relying solely on its binary threshold mechanism to control optimization.
+Impact of SFT Regularization.
+Our experimental results reveal a consistent trend: as $\lambda$ increases from 0.0 to 5.0, performance on AlpacaEval2 LC steadily declines from 34.1% to 27.8%.
+This progressive degradation suggests that stronger regularization toward the initial SFT model actually hinders effective preference learning in this context.
+The optimal performance occurs at $\lambda=0.0$, which effectively transforms SLiC-HF into a variant of RePO.
+Different Optimization Challenges.
+These findings suggest that SFT regularization and RePO’s threshold-based filtering address fundamentally different optimization challenges.
+While SFT regularization was originally introduced to mitigate catastrophic forgetting and preserve general capabilities, our results indicate that for direct preference optimization, such regularization is unnecessary and potentially counterproductive.
+Instead, RePO’s selective gradient application through its threshold mechanism appears sufficient to prevent over-optimization while maintaining effective preference learning.
+This analysis complements our main findings and further supports our hypothesis that carefully designed filtering mechanisms can effectively replace more complex regularization schemes in preference optimization.
+Table 6: The hyperparameter $\lambda$ in SliC-HF used for each Llama3-Instruct v0.2.
+SLiC-HF $\lambda=0.0$ $\lambda=0.1$ $\lambda=0.5$ $\lambda=1.0$ $\lambda=3.0$ $\lambda=5.0$
+AlpacaEval2 LC 34.1 33.9 32.8 30.8 28.6 27.8
+D.6 Detailed Analysis of the Relationship Between DPO and RePO
+To further investigate the relationship between DPO and RePO, we conducted additional experiments examining how different formulation components affect performance.
+As shown in Table 7, we systematically evaluated five variants that decompose the key elements of each approach.
+DPO as a Special Case of SimPO.
+From a mathematical perspective, DPO can be viewed as a specific instance of SimPO where $\gamma=\log\pi_{\text{ref}}(y_{w}\mid x)-\log\pi_{\text{ref}}(y_{l}\mid x)$ (ignoring length normalization for equivalence).
+This connection highlights how DPO implicitly defines its target margin based on reference model probabilities rather than using an explicit hyperparameter.
+Impact of ReLU Without Explicit Margin.
+The second row of Table 7 shows that directly replacing log-sigmoid with ReLU while maintaining DPO’s implicit margin definition leads to catastrophic performance degradation (-44% on AlpacaEval LC, -47% on AlpacaEval WR, and -26% on Arena-Hard).
+This dramatic decline reveals that the binary threshold mechanism of ReLU is only effective when paired with an appropriate explicit margin parameter.
+The Critical Role of $\gamma$.
+Rows 3-5 demonstrate that adding an explicit $\gamma$ parameter consistently improves performance across all metrics regardless of whether log-sigmoid, ReLU, or their combination is used.
+The most substantial gains appear when ReLU and $\gamma$ are combined (+4.4% LC, +9.1% WR on AlpacaEval), supporting our hypothesis that explicit threshold-based filtering effectively controls over-optimization.
+Complementary Mechanisms.
+Interestingly, the combination of both mechanisms (row 5) yields the highest overall performance, suggesting that while RePO’s binary filtering mechanism addresses the core over-optimization challenge, the continuous weighting from log-sigmoid may provide complementary benefits for fine-grained preference learning.
+Table 7: Impact Analysis of $\gamma$ Scaling and ReLU Mechanisms in DPO Training.
+Benchmark results on AlpacaEval 2.0 (AE2) and Arena-Hard (AH) demonstrate percentage point changes in Length-Controlled Win Rate (LC-WR) and Base Win Rate (WR) for Llama3-Instruct-v0.2 (8B).
+Values represent relative performance deltas (%) compared to standard DPO baseline.
+Method $\gamma$ ReLU $\log\sigma$ Llama3-Instruct v0.2 (8B)
+AE 2 AH
+LC WR WR
+$\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ 48.2 47.5 35.2
+$\text{ReLU}\left(-\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right)$ $26.9^{\color[rgb]{0,0,0}-44\%}$ $25.1^{\color[rgb]{0,0,0}-47\%}$ $26.2^{\color[rgb]{0,0,0}-26\%}$
+$\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\gamma\right)$ $50.0^{\color[rgb]{0,0,0}+3.7\%}$ $50.7^{\color[rgb]{0,0,0}+6.7\%}$ $36.8^{\color[rgb]{0,0,0}+4.5\%}$
+$\text{ReLU}\left(-\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\gamma\right)\right)$ $50.3^{\color[rgb]{0,0,0}+4.4\%}$ $51.8^{\color[rgb]{0,0,0}+9.1\%}$ $38.2^{\color[rgb]{0,0,0}+8.5\%}$
+$\log\sigma\left(-\text{ReLU}\left(-\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\gamma\right)\right)\right)$ $50.8^{\color[rgb]{0,0,0}+5.4\%}$ $52.2^{\color[rgb]{0,0,0}+9.9\%}$ $37.2^{\color[rgb]{0,0,0}+5.7\%}$
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: see abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: see Section 6.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: see Appendix C.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: see Appendix D.1.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  25. 6.
+    Experimental setting/details
+  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  27. Answer: [Yes]
+  28. Justification: see Appendix D.1.
+  29. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  30. 7.
+    Experiment statistical significance
+  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  32. Answer: [Yes]
+  33. Justification: The results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+  34. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  35. 8.
+    Experiments compute resources
+  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  37. Answer: [Yes]
+  38. Justification: We carried out all computational tasks on a suite of four 80GB A100 GPUs.
+  39. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  40. 9.
+    Code of ethics
+  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  42. Answer: [Yes]
+  43. Justification: Our research has been conducted with strict adherence to the NeurIPS Code of Ethics.
+  44. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  45. 10.
+    Broader impacts
+  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  47. Answer: [Yes]
+  48. Justification: see Appendix B.
+  49. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  50. 11.
+    Safeguards
+  51.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  52. Answer: [N/A]
+  53. Justification: The paper poses no such risks.
+  54. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  55. 12.
+    Licenses for existing assets
+  56.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  57. Answer: [N/A]
+  58. Justification: The paper does not use existing assets.
+  59. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  60. 13.
+    New assets
+  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  62. Answer: [N/A]
+  63. Justification: The paper does not release new assets.
+  64. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  65. 14.
+    Crowdsourcing and research with human subjects
+  66.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  67. Answer: [N/A]
+  68. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  69. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  70. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  71.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  72. Answer: [N/A]
+  73. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  74. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  75. 16.
+    Declaration of LLM usage
+  76.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  77. Answer: [N/A]
+  78. Justification: The LLM is used only for writing, editing.
+  79. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.